### v1

# Polytope Volume Monitoring Problem: Formulation and Solution via Parametric Linear Program Based Control Barrier Function

Abstract
Motivated by the latest research on feasible space monitoring of multiple control barrier functions (CBFs) as well as polytopic collision avoidance, this paper studies the Polytope Volume Monitoring (PVM) problem, whose goal is to design a control law for inputs of nonlinear systems to prevent the volume of some state-dependent polytope from decreasing to zero. Recent studies have explored the idea of applying Chebyshev ball method in optimization theory to solve the case study of PVM; however, the underlying difficulties caused by nonsmoothness have not been addressed. This paper continues the study on this topic, where our main contribution is to establish the relationship between nonsmooth CBF and parametric optimization theory through directional derivatives for the first time, so as to solve PVM problems more conveniently. In detail, inspired by Chebyshev ball approach, a parametric linear program (PLP) based nonsmooth barrier function candidate is established for PVM, and then, sufficient conditions for it to be a nonsmooth CBF are proposed, based on which a quadratic program (QP) based safety filter with guaranteed feasibility is proposed to address PVM problems. Finally, a numerical simulation example is given to show the efficiency of the proposed safety filter.
I Introduction
The control barrier function (CBF) based control design method is popular for safety-critical systems[1]. CBF-based controllers are typically implemented pointwisely in the form of quadratic program (QP), where CBF conditions and input constraints are incorporated as linear inequalities. Due to its advantages of low computational complexity and formal safety guarantees, CBF methods are successfully applied in various autonomous robotic systems [2, 3, 4].
The appearance of multiple CBFs and input constraints would impose great challenges in the feasibility of the resulting QP-based control laws, which has been paid much attention recently [5, 6, 7, 8, 9]. Some primary works study sufficient conditions of the compatibility of multiple CBFs [6, 7]. More recently, in [8], the nonsmooth CBF, dating backing to [5], is introduced as a systematic tool to address challenges in guaranteeing the feasibility and continuity. Meanwhile, a simpler way of addressing the feasibility problem appears in [9]. In detail, “Feasible CBF Space” is defined to describe the state-dependant polytope (in input space) formulated by CBF conditions and input constraints. Then, Feasible Space Monitoring (FSM) problem is established to monitor the feasibility of the resulting QP-based control laws, for which Chebyshev ball and inscribing ellipsoid methods are proposed to build parametric linear program (PLP) based CBFs. For engineering implementations, the differentiable optimization toolbox is applied to approximate the gradient of the resulting (nonsmooth implicit) barrier function candidate by some heuristic gradient values. As concluded in [9], the nonsmooth control theoretic considerations are left for future work. To the best of our knowledge, this problem has been open until now. Such observation motivates this paper.
Meanwhile, researchers have been gradually interested in building CBFs via parametric optimization-defined metrics for related problems, such as network connectivity maintenance [10] and polytopic collision avoidance (PCA) [11, 12, 13, 14, 15, 16]. However, it is difficult to extend these results to solve the above open problem in FSM. Roughly speaking, there are three potential ways, yet all meet significant difficulties: 1) In [10], the Laplacian’s eigenvalue is chosen as a barrier function candidate, then, a nonsmooth CBF condition in the generalized gradient form is formulated. The idea of building nonsmooth CBF conditions based on generalized gradient can only be applied to specific problems, and it is not applicable to other parametric optimization problems. 2) In [15, 16], the minimum distance field for PCA, i.e., a parametric quadratic program defined metric, is used to formulate a nonsmooth CBF condition based on the directional derivative of the KKT solution. How to extend it to other problems is also not obvious. 3) Recently, the nonsmoothness in PCA has been avoided by circumscribing polytopes into strictly convex shapes to construct differentiable optimization defined CBFs [11, 12, 13] or by proposing some nonconservative approximation of signed distance field to establish the optimization-free smooth CBF [14]. This indirect approach avoids the challenges of nonsmooth parametric optimization, transforming the original simple LP into a nonlinear program. In summary, the existing works about parametric optimization defined CBF methods [10, 11, 12, 13, 14, 15, 16] could not be directly applied to address the difficulties of nonsmoothness in FSM problems.
This paper extends the study of Feasible Space Monitoring [9] and addresses the unresolved difficulties in nonsmooth parametric optimization based CBFs [10, 11, 12, 13, 14, 15, 16]. First, a broader class of control problems is formulated, called Polytope Volume Monitoring (PVM). By identifying this kind of problems, FSM problem [9] and Polytopic Intersectability Maintenance (PIM), i.e., the opposite problem of PCA [11, 12, 13, 14, 15], can be regarded as different special examples of PVM, as seen in Fig. 1. This generalization enhances the applicability of the proposed method to a wider range of problems. Second, to address difficulties in analysis and synthesis caused by inherent nonsmoothness, this paper attempts to establish a bridge between the directional derivative defined nonsmooth CBF [17] and parametric optimization theory [18]. In detail, our main contributions can be summarized as follows:
  • 1)
    Compared with existing works [10, 11, 12, 13, 14, 15, 16], this paper gives a direct way to analyze and synthesize parametric optimization defined nonsmooth CBFs by merging [17] and [18] into a unified framework for the first time.
  • 2)
    Within this unified framework, a parametric LP based nonsmooth barrier function candidate is established for PVM, based on which sufficient conditions of being a nonsmooth CBF are proposed. Then a slacked QP-based safety filter with guaranteed feasibility is proposed to solve the PVM problem. The above nonsmooth analysis and considerations are not addressed in [9].
  • 3)
    The efficiency of the proposed safety filter is validated in a numerical simulation where a vehicle is driven to perform safe navigation, i.e., a reach-avoid task under multiple obstacles and input constraints.
The paper is organized as follows. The CBF method is briefly reviewed in Section II, and problem formulation and preliminary analysis are given in Section III. Section IV reports the main results. Numerical simulation is performed in Section V. Appendix includes the referred nonsmooth parametric optimization theory and proofs.
Notations: Throughout this paper, $\operatorname{col}(x_{1},x_{2}):=[x_{1}^{\top},x_{2}^{\top}]^{\top}$ for any two column vectors $x_{1},x_{2}$. For any vector $x$, $\|x\|$ denotes its Euclidean norm. $|\mathcal{J}|$ denotes the cardinality of any finite number set $\mathcal{J}$. 00 and $\emptyset$ denote the zero vector and empty set respectively in suitable dimension. The interior and boundary of a set $\mathcal{C}$ are indicated by $\operatorname{Int}(\mathcal{C})$ and $\partial\mathcal{C}$. A continuous function $\alpha:\mathbb{R}\rightarrow\mathbb{R}$ is called in extended class-$\mathcal{K}$ if it is strictly increasing and with $\alpha(0)=0$ [1]. For any function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$, if exists, its generalized gradient at $x$ is denoted by $\partial f(x)$, and $f^{\prime}(x;d)$ denotes its directional derivative at $x$ along direction $d$. For brevity, its partial derivative $\frac{\partial f}{\partial x}(x)$, expressed as a row vector, is denoted by $\partial_{x}f$ sometimes.
II Preliminaries: Control Barrier Function
Consider a control-affine nonliear system
$\dot{x}(t)=f_{0}(x(t),u(t)):=f(x(t))+g(x(t))u(t),$ (1)
where $x\in\mathbb{R}^{n}$ is state, $u\in\mathcal{D}_{u}\subset\mathbb{R}^{m}$ is control input, and $f,g$ are Lipschitz continuous. Next, the nonsmooth barrier function (NBF) candidate [5] is introduced.
Definition 1.
Given a nonempty closed set $\mathcal{C}\subset\mathbb{R}^{n}$ with no isolated point, a locally Lipschitz continuous function $h$ : $\mathbb{R}^{n}\rightarrow\mathbb{R}$ is called a NBF candidate on set $\mathcal{C}$ if
$h(x)$ $=0,~{}\forall x\in\partial\mathcal{C};\quad h(x)>0,~{}\forall x\in\operatorname{Int}(\mathcal{C});$ (2)
$h(x)$ $<0,~{}\forall x\in\mathbb{R}^{n}/\mathcal{C}.$
Function $h$ is called a differentiable CBF to (1) if BF candidate $h$ is further differentiable and there exists an enxtended $\mathcal{K}$-class function $\alpha$ such that for all $x$ it holds $\sup_{u\in\mathcal{D}_{u}}\left\{\frac{\partial h}{\partial x}(x)f_{0}(x,u)\right\}\geq-\alpha(h(x)).$ For the demands of this paper, we only require from now on that $h$ is Lipschitz continuous and the (one-sided) directional derivative of $h(x)$ at point $x$ along the direction $\dot{x}$ exists, i.e., the following limit is finite and abbreviated as
$h^{\prime}(x(t);\dot{x}(t))$ $:=\lim_{\sigma\rightarrow 0^{+}}\frac{h(x(t)+\sigma\dot{x}(t))-h(x(t))}{\sigma}.$ (3)
By doing this, one can relax the differentiability of $h$ and define a broader class of CBFs as follows.
Definition 2.
A locally Lipschitz continuous function $h:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is a nonsmooth control barrier function (NCBF) for system (1) if $h$ is a NBF candidate on the closed set $\mathcal{C}\subset\mathbb{R}^{n}$, and there exists an extended class-$\mathcal{K}$ function $\alpha:\mathbb{R}\rightarrow\mathbb{R}$, an open set $\mathcal{D}$ that satisfies $\mathcal{C}\subset\mathcal{D}\subseteq\mathbb{R}^{n}$, such that:
$\sup_{u\in\mathcal{D}_{u}}\left\{h^{\prime}\left(x;f_{0}(x,u)\right)\right\}\geq-\alpha(h(x)),\quad\forall x\in\mathcal{D}.$ (4)
In [17], the same limitation value as $h^{\prime}(x;f_{0}(x,u))$ is used to define NCBF, yet without introducing directional derivative formally. In this paper, the directional derivative defined NCBF (4) is introduced to build a bridge between the NCBF theory and the parametric optimization theory.
Lemma 1.
[17, Theorem 1]  If $h$ is a NCBF for system (1), then the closed set $\mathcal{C}$ is forward control-invariant, that is, there exists $u(t)\in\mathcal{D}_{u}$, such that $x_{0}\in\mathcal{C}$ implies that $x(t)\in\mathcal{C},~{}\forall t\in\left[0,T\right]$, for every solution.
When $h(x)$ is differentiable, it holds that $h^{\prime}(x;f_{0}(x,u))=\frac{\partial h}{\partial x}(x)\left(f(x)+g(x)u\right)$, which is a linear function of input $u$ at every $x$, and the following is a QP-defined control law:
$u^{*}(x)=\operatorname*{argmin}_{u\in\mathcal{D}_{u}}~{}$ $\|u-u_{0}\|^{2},$ (5a)
s.t. $h^{\prime}\left(x;f_{0}(x,u)\right)\geq-\alpha(h(x)),$ (5b)
where $u_{0}$ is any given continuous nominal control law.
When $h(x)$ is no longer differentiable but directional derivative $h^{\prime}(x;f_{0}(x,u))$ exists, the inequality (5b) may no longer be a linear constraint of input $u$.
Remark 1.
Different from (4), the more popular NCBF is defined by the set-valued Lie derivative in [5], [19, 10]:
$\sup_{u\in\mathcal{D}_{u}}\left\{\min_{\zeta\in\partial h(x)}\zeta^{\top}f_{0}(x,u)\right\}\geq-\alpha(h(x)),\quad\forall x\in\mathcal{D},$ (6)
[…]
III Problem Formulation and Analysis
III-A Problem Statement
For simplicity of discussions, consider the control-affine system (1) with a polytopic set-valued output and a polytopic input constraint:
$\dot{x}$ $=f(x)+g(x)u,$ (7a)
$y$ $=\Phi(x):=\left\{c\in\mathbb{R}^{l}\mid a_{j}(x)^{\top}c\leq b_{j}(x),j\in\mathcal{J}\right\},$ (7b)
$u$ $\in\Psi(x):=\left\{u\in\mathbb{R}^{m}\mid\mathfrak{a}_{\mathfrak{i}}(x)^{\top}u\leq\mathfrak{b}_{\mathfrak{i}}(x),\mathfrak{i}\in\mathfrak{J}\right\},$ (7c)
where $\mathcal{J}:=\{1,2,...,N\}$ and $\mathfrak{J}:=\{1,2,...,\mathfrak{N}\}$. For convenience, the functions appearing in definitions of $\Phi(x)$ and $\Psi(x)$ are assumed to be continuously differentiable, and the following assumptions are imposed to develop an implementable controller.
Assumption 1.
The state-dependent polytopic sets $\Phi(x)$, $\Psi(x)$ are both nonempty and uniformly bounded for all $x$, i.e., there exist compact sets $\mathcal{D}_{y}\subset\mathbb{R}^{l}$ and $\mathcal{D}_{u}\subset\mathbb{R}^{m}$, such that $\Phi(x)\subseteq\mathcal{D}_{y}$ and $\Psi(x)\subseteq\mathcal{D}_{u}$ respectively. Additionally, vectors $a_{j}(x),j\in\mathcal{J}$ are nonzero at any $x$.
Without loss of generality, Assumption 1 is made for some technical considerations. Especially, the non-emptiness of $\Phi(x)$ is essential to guarantee that some metric is well-defined in the following sections. In fact, this paper will attempt to synthesize a controller to prevent the interior of $\Phi(x(t))$ from vanishing, as formally stated below.
Figure 1: Illustration of the PVM problem and its examples. (a) Problem statement and preliminary analysis. (b) Example 1 in input space. (c) Example 2 in state space.
Polytope Volume Monitoring Problem: Consider the system (7) with a polytopic set-valued output $\Phi(x)$ and a polytopic input constraint set $\Psi(x)$. The goal is to design a control law for $u\in\Psi(x)$ to prevent the volume of the polytope $\Phi(x(t))$ from being less than a pre-defined positive value, i.e.,
$\mathcal{V}(x(t)):=\int_{\Phi(x(t))}\mathrm{d}s\geq\epsilon,~{}\forall t\geq 0,$ (8)
where $\epsilon>0$ is any user-defined threshold value.
There are some typical situations where the volume of state-dependant polytope $\Phi(x(t))$ might be zero if the input $u$ is not designed properly.
Example 1 (Feasible Space Monitoring [9]).
In Fig. 1(b), given the polytopic input constraint set $\mathcal{U}$, by constructing the set composed of multiple differentiable CBFs $h_{i},i\in\mathcal{I}$: $\mathcal{U}_{c}(x):=\{u\mid\frac{\partial h_{i}}{\partial x}(x)\left(f(x)+g(x)u\right)\geq-\alpha_{i}(h_{i}(x)),i\in\mathcal{I}\}$, the FSM problem aims to design $u$ such that
$\mathcal{U}\cap\mathcal{U}_{c}(x(t))\neq\emptyset,~{}\forall t.$ (9)
By choosing $\Phi(x)=\Psi(x)=\mathcal{U}\cap\mathcal{U}_{c}(x)$, feasible space monitoring of any QP-based controllers can be regarded as case studies of the formulated PVM problem.
In addition to the above examples in input space, PVM problems also appear in state space. In many applications, it is necessary to maintain some kind of connectivity of two sets, for instance, a person standing on vehicle $i$ needs to transfer to vehicle $j$. During this process, $\mathcal{P}^{i}(x^{i})$ and $\mathcal{P}^{j}(x^{j})$ should be connected, as shown in (10).
Example 2 (Polytopic Intersectability Maintenance).
Consider two standard nonholonomic vehicles $i,j$, for each $\mathfrak{i}\in\{i,j\}$, $x^{\mathfrak{i}}=[{p}_{x}^{\mathfrak{i}},{p}_{y}^{\mathfrak{i}},{\theta}^{\mathfrak{i}}]^{\top}$ and $u^{\mathfrak{i}}=[v^{\mathfrak{i}},\omega^{\mathfrak{i}}]^{\top}\in\mathcal{U}^{\mathfrak{i}}$ are state and input vectors respectively, where $\mathcal{U}^{\mathfrak{i}}$ denotes a hyperrectangle. Let sets $\mathcal{P}^{i}(x^{i})$ and $\mathcal{P}^{j}(x^{j})$ represent their physical domain or sensor ranges respectively, as shown in Fig. 1(c). The goal of polytopic intersectability maintenance is to maintain the intersection of two polytopes nonempty:
$\mathcal{P}^{i}(x^{i}(t))\cap\mathcal{P}^{j}(x^{j}(t))\neq\emptyset,~{}\forall t,$ (10)
which is opposite to the goals of polytopic collision avoidance [11, 12, 13, 14, 15, 16]. From the perspective of PVM, $\Phi(x)=\mathcal{P}^{i}(x^{i})\cap\mathcal{P}^{j}(x^{j})$ and $\Psi(x)=\mathcal{U}^{i}\times\mathcal{U}^{j}$ in this example.
III-B Preliminary Analysis
The center $c\in\mathbb{R}^{l}$ and radius $r\geq 0$ of the largest sphere contained inside the polytope $\Phi(x)$ defined in (7b), called Chebyshev ball [21, Chapter 8], can be computed by the following Linear Program at every $x$:
$r^{*}(x):=$ $\max_{c,r}~{}r=-\min_{c,r}~{}-r,$ (11a)
s.t. $a_{j}^{\top}(x)c+\|a_{j}(x)\|r\leq b_{j}(x),~{}\forall j\in\mathcal{J},$ (11b)
then from the geometric meaning in Fig. 1(a), it holds that
$\mathcal{V}(x(t))\geq\pi(r^{*}(x(t)))^{2}\geq\epsilon,$ (12)
from which one knows the control goal in (8) would hold if
$r^{*}(x(t))\geq\epsilon_{0}:=\sqrt{\frac{\epsilon}{\pi}},~{}\forall t\geq 0.$ (13)
Remark 2.
[…]
IV Main results
To solve the aforementioned open problem, the nonsmooth parametric optimization theory, as shown in Appendix -A, is introduced to develop a general approach for synthesizing a NCBF-based safety filter to solve PVM problems for the first time. Without calculating the generalized gradient $\partial r^{*}(x)$, the NCBF condition is established based on its directional derivative for convenience.
IV-A Establishing a Parameteric LP-defined NBF Candidate
Inspired by (13), one can define an implicit function $h_{r}(x)$ via the following parametric linear program:
$h_{r}(x):=$ $r^{*}(x)-\epsilon_{0}=-\epsilon_{0}-\min_{z\in F_{P}(x)}~{}q^{\top}z,$ (14)
where $q:=\operatorname{col}{(0_{l\times 1},-1)}$ and $z:=\operatorname{col}{(c,r)}\in\mathbb{R}^{l+1}$. The feasible set $F_{P}(x)$ is defined as
$F_{P}(x):=\left\{z\mid\tilde{\mathbf{A}}(x)z\leq\tilde{\mathbf{b}}(x)\right\},$ (15)
where $\tilde{\mathbf{A}}(x)\in\mathbb{R}^{N\times(l+1)}$ and $\tilde{\mathbf{b}}(x)\in\mathbb{R}^{N}$ are
$\tilde{\mathbf{A}}(x):=\left[\begin{array}[]{cc}\tilde{a}_{1}(x)^{\top}\\ \tilde{a}_{2}(x)^{\top}\\ \cdots\cdots\\ \tilde{a}_{N}(x)^{\top}\end{array}\right],\quad\tilde{\mathbf{b}}(x):=\left[\begin{array}[]{c}b_{1}(x)\\ b_{2}(x)\\ \cdots\cdots\\ b_{N}(x)\end{array}\right],$ (24)
with $\tilde{a}_{j}(x):=\operatorname{col}(a_{j}(x),\|a_{j}(x)\|)$ for every $j=1,2,...,N$.
Next, preliminary properties of the PLP in (14) are shown first, where Mangasarian Fromovitz Constraint Qualification (MFCQ), as in Definition 3, is a key condition in studying the continuity and directional differentiability of $r^{*}(x)$.
Proposition 1.
For the value function $h_{r}(x)$ of the PLP in (14), the following conclusions hold:
  • i)
    The feasible set $F_{P}(x)$ is nonempty for every $x$ and uniformly bounded for all $x$.
  • ii)
    For every $x$, $h_{r}(x)$ is finite, and the solution map
    $S(x):=$ $\operatorname*{argmin}_{z\in F_{P}(x)}~{}q^{\top}z,$ (25)
    is on the boundary of $F_{P}(x)$, i.e., $S(x)\subset\partial F_{P}(x)$.
  • iii)
    For every $x$, the MFCQ holds at any $z\in F_{P}(x)$.
Proof.
To avoid breaking the flow of the exposition, please see the note in Appendix -B. ∎
Based on Proposition 1, one can prove that $h_{r}(x)$ can be chosen as a NBF candidate in the sense of (2).
Theorem 1.
The Parametric LP defined implicit function $h_{r}(x)$ in (14) is locally Lipschitz continuous, and it is a NBF candidate on closed set $\mathcal{S}:=\{x:h_{r}(x)\geq 0\}$.
Proof.
Recalling Proposition 1 i) and iii) that ${F}_{P}(x)$ is uniformly bounded for all $x$ and the MFCQ condition holds, according to Lemma 2, $h_{r}(x)$ is locally Lipschitz continuous at every $x$. From (14), the conditions in (2) hold for $h_{r}(x)$, hence $h_{r}(x)$ is a NBF candidate on $\mathcal{S}$. ∎
Before going to develop the theoretical results, the geometric structure of the solution map $S(x)$ is discussed, which is helpful to understand the expression of $h_{r}^{\prime}(x;\dot{x})$ in (28). First, from the definition (25), $S(x)$ can be rewritten as
$S(x)$ $=T(x)\times\left\{r\mid r=r^{*}(x)\right\},$
where
$T(x):=\bigcap_{j\in\mathcal{J}}\left\{c\mid a_{j}^{\top}(x)c\leq b_{j}(x)-\|a_{j}(x)\|r^{*}(x)\right\}.$
Since $r^{*}(x)\geq 0$, one has
$T(x)\subseteq\left\{c\mid a_{j}^{\top}(x)c\leq b_{j}(x),\forall j\in\mathcal{J}\right\}=\Phi(x)\subset\mathcal{D}_{y}.$
Because $T(x)$ is expressed as the intersection of halfspcaes, it is still a (convex and compact) polytope in $\mathcal{D}_{y}\subset\mathbb{R}^{l}$, whose dimension, lower than $\Phi(x)$, may be $10,1,2,...,l-$ 1. In the above, dimension 00 means that $T(x)$ reduces to a singleton.
IV-B Building the NCBF Condition
This section aims to establish the NCBF condition based on existing results about the directional differentiability of parametric convex programs in Lemma 3.
In detail, recalling the parameteric linear program in (14):
$v_{r}(x):=\min_{z\in F_{P}(x)}~{}q^{\top}z=-h_{r}(x)-\epsilon_{0},$
then the directional derivatives of $v_{r}(x)$ and $h_{r}(x)$ are studied.
First, according to the feasible set $F_{P}(\cdot)$ in (15), the active index set [^1] Inspired by [19], $\mathcal{J}_{0}({z},x)$ can be replaced by the socalled almost-active index set $\mathcal{J}_{\varepsilon}({z},x):=\{j\in\mathcal{J}\mid\|\tilde{a}_{j}(x)^{\top}{z}-b_{j}(x)\|\leq\varepsilon\}$ for real implementation, where $\varepsilon$ is a sufficiently small scalar, such as $\varepsilon=10^{-4}$. is defined for every ${x}$ and feasible ${z}\in F_{P}({x})$ as:
$\mathcal{J}_{0}({z},x):=$$\left\{j\in\mathcal{J}\mid\tilde{a}_{j}(x)^{\top}{z}=b_{j}(x)\right\}$ (26)
$=$ $\left\{j\in\mathcal{J}\mid a_{j}^{\top}(x)c+\|a_{j}(x)\|r=b_{j}(x)\right\},$
where $\tilde{a}_{j}(x):=\operatorname{col}(a_{j}(x),\|a_{j}(x)\|)$.
For every $\bar{z},\bar{x}$ satisfying $\bar{z}\in S(\bar{x})$, recalling $S(\bar{x})\subset\partial F_{P}(\bar{x})$, it holds that
$\mathcal{J}_{0}(\bar{z},\bar{x})\neq\emptyset,\text{ i.e., }|\mathcal{J}_{0}(\bar{z},\bar{x})|\geq 1.$
Based on (24), one can further define the active matrix of $\tilde{\mathbf{A}}(x)$ and the active vector of $\tilde{\mathbf{b}}(x)$ for every $x$ as:
$\tilde{\mathbf{A}}_{\mathcal{J}_{0}(\bar{z},\bar{x})}(x):=\left[\begin{array}[]{c}\tilde{a}_{j_{1}}(x)^{\top}\\ \tilde{a}_{j_{2}}(x)^{\top}\\ ......\\ \tilde{a}_{j_{m}}(x)^{\top}\end{array}\right],\tilde{\mathbf{b}}_{\mathcal{J}_{0}(\bar{z},\bar{x})}(x):=\left[\begin{array}[]{c}b_{j_{1}}(x)\\ b_{j_{2}}(x)\\ ......\\ b_{j_{m}}(x)\end{array}\right],$ (27)
where $j_{1},j_{2},...,j_{m}$ represent all indices in $\mathcal{J}_{0}(\bar{z},\bar{x})$. In the above, variables “$\bar{z},\bar{x}$”, appearing in the active set $\mathcal{J}_{0}(\cdot,\cdot)$, are introduced to make a difference from the variable “$x$” in $\tilde{a}_{j}(x)$ and $b_{j}(x)$. Then the directional derivative of $h_{r}(x)$ along the system can be obtained through Lemma 3.
Proposition 2.
The directional derivative of $h_{r}(x)$ in (14) at $x$ along $\dot{x}$ can be expressed as
$h^{\prime}_{r}(x;\dot{x})$ $=\max_{z\in S(x)}\min_{\mu\in M(z,x)}-\partial_{x}L(\mu,z,x)\dot{x},$ (28)
where the solution map $S(x)$ is defined as in (25),
$\partial_{x}L(\mu,z,x)$ $=\sum_{j\in\mathcal{J}_{0}({z},{x})}\mu_{j}\left(z^{\top}\frac{\partial\tilde{{a}}_{j}}{\partial x}(x)-\frac{\partial{{b}}_{j}}{\partial x}(x)\right),$ (29)
and the dual feasible set
$M(z,x)=$$\left\{\mu\mid\tilde{\mathbf{A}}^{\top}_{\mathcal{J}_{0}({z},{x})}(x)\mu=-q,~{}\mu\geq 0\right\},$ (30)
is bounded (compact).
Proof.
To avoid breaking the flow of the exposition, please see the note in Appendix -B. ∎
Based on Proposition 2 and Definition 2, the NCBF condition to guaranteeing $h_{r}(x(t))\geq 0,~{}\forall t$ along the system (7a) can be briefly expressed as follows:
$\max_{\begin{subarray}{c}z\in S(x)\\ u\in\Psi(x)\end{subarray}}\min_{\mu\in M(z,x)}-\partial_{x}L(\mu,z,x)\left(f+gu\right)\geq-\alpha(h_{r}(x)).$ (31)
It is worth mentioning that condition (31) is more complex than the regular NCBF condition (6). Due to the (double-layer) max-min optimization form, it is not obvious to calculate the generalized gradient of $h_{r}(x)$. To simplify (31) into a more practical form that is convenient to synthesize control laws, a sufficient condition is established.
Theorem 2.
The BF candidate $h_{r}(\cdot)$ is further a NCBF on $\mathcal{S}$, i.e., the NCBF condition in (4) holds for $h_{r}(\cdot)$, if there exist $\alpha,\epsilon>0$, for all $x\in\mathcal{D}_{\epsilon}$, $\exists u\in\Psi(x)$, such that
$\min_{\mu\in M(z^{*}(x),x)}-\partial_{x}L(\mu,z^{*}(x),x)\left(f+gu\right)\geq-\alpha h_{r},$ (32)
where $z^{*}(x)=\operatorname{col}{(c^{*}(x),r^{*}(x))}$ is chosen from $S(x)$ and the expression of $\partial_{x}L(\cdot)$ is given in (29).
Proof.
Based on (28) and (32), one has
$h^{\prime}_{r}(x;f(x)+g(x)u)$
$=\max_{z\in S(x)}\min_{\mu\in M(z,x)}-\partial_{x}L(\mu,z,x)\left(f+gu\right)$
$\geq~{}\min_{\mu\in M(z^{*}(x),x)}-\partial_{x}L(\mu,z^{*}(x),x)\left(f+gu\right)$
$\geq-\alpha h_{r},$
which indicates the NCBF condition (31) holds. ∎
Remark 3.
[…]
IV-C Synthesizing a NCBF-based Safety Filter
Based on Theorem 2, this section attempts to synthesize a NCBF-based safety filter to solve PVM problems.
Given the chosen point $z^{*}(x)\in S(x)$, recalling the expression in (30), the dual set $M(z^{*}(x),x)$ is a bounded polytope in $\mathbb{R}^{|\mathcal{J}_{0}(z^{*}(x),x)|}$. From the famous Minkowski-Weyl’s Theorem [23] that every polytope can be represented in two equivalent ways (i.e., halfspace- and vertex- representations), there exists a finite point set $V_{M}(x)=\left\{{\mu}^{k}(x)\mid k\in\mathcal{I}_{V_{M}}(x)\right\}$, called vertexes [^2] Till now, it is convenient to obtain the vertex rep. $V_{M}(x)$ from half-space rep. of $M(z^{*}(x),x)$ in (30) automatically, such as the open-source toolbox $\mathrm{pypoman}$ [24]., such that
$M(z^{*}(x),x)=\operatorname{co}\left(\{{\mu}^{k}(x)\mid k\in\mathcal{I}_{V_{M}}(x)\}\right),$ (33)
where $\operatorname{co}(\cdot)$ denotes the convex hull operator of a set, and $\mathcal{I}_{V_{M}}(x)$ is the index set of ${V_{M}}(x)$.
Based on (33), a direct corollary of Theorem 2 is shown.
Corollary 1.
The BF candidate $h_{r}(\cdot)$ is further a NCBF on $\mathcal{S}$, i.e., the NCBF condition in (4) holds for $h_{r}(\cdot)$, if there exist $\alpha,\epsilon>0$, for all $x\in\mathcal{D}_{\epsilon}$, $\exists u\in\Psi(x)$, such that
$\Gamma(\mu^{k}(x),x,u)\geq-\alpha h_{r}(x),~{}\forall k\in\mathcal{I}_{V_{M}}(x),$ (34)
where $\Gamma(\mu,x,u):=-\partial_{x}L(\mu,z^{*}(x),x)\left(f+gu\right)$ for any $\mu$.
Proof.
From the expression of $\partial_{x}L(\cdot)$ in (29), one can know that the scalar function $\partial_{x}L(\mu,z^{*},x)\left(f+gu\right)$ is a linear function of $\mu,z^{*},u$ respectively, hence the minimization optimization about $\mu$ in (32) is a LP at every $x$. Recalling the fact in [25, Sec 2.5, Theorem and Corollary 2] that the minimum value of a LP is obtained at the vertexes of the polytope, then
$\min_{\mu\in M(z^{*}(x),x)}\Gamma(\mu,x,u)$ $=\min_{\mu\in V_{M}(x)}\Gamma(\mu,x,u)$
$=\min_{{k}\in\mathcal{I}_{V_{M}}(x)}\Gamma(\mu^{k},x,u),$
which can be further combined with (32) to verify (34). ∎
Based on Corollary 1, a QP-based safety filter can be further built to solve Polytope Volume Monitoring:
$\bar{u}^{*}(x)=\operatorname*{argmin}_{u,\delta}~{}(u-u_{0})^{\top}Q(u-u_{0})+\gamma\delta^{2},$ (35a)
[…]
$\quad\quad u\in\Psi(x),~{}\delta\geq 0,$ (35c)
where $\bar{u}^{*}(x):=\operatorname{col}(u^{*}(x),\delta^{*}(x))$, $u_{0}$ is any given continuous nominal control law, and the slack variable $\delta$ is introduced to guarantee the feasibility of such QP. $Q\in\mathbb{R}^{m\times m},\gamma>0$ are the (positive-definite) weighting matrix and scalar. The structure of the proposed safety filter (35) is summarized in Fig. 2.
Figure 2: A summary of the structure of the proposed safety filter (35).
Till now, the final theorem is shown to conclude this paper.
Theorem 3.
Consider the designed safety filter in the form of (35). For any system trajectory $x(t)$ starting from $x_{0}\in\mathcal{S}=\{x:h_{r}(x)\geq 0\}$, if
$u^{*}(x(t))\text{ is continuous, and }\delta^{*}(x(t))\equiv 0,$ (36)
then $x(t)\in\mathcal{S},\forall t\geq 0$, which implies that the PVM problem is solved. As a consequence,
$\Phi(x(t))\neq\emptyset,~{}u^{*}(x(t))\in\Psi(x(t)),\quad\forall t\geq 0,$ (37)
i.e., the output set-valued mapping never be empty and meanwhile the input constraint is never violated.
Proof.
Since Eq. (35b) denotes the soft constraints, the feasibility of the QP-based safety filter is guaranteed. The solution of closed-loop system is well-defined by the assumption that $u^{*}(x(t))$ is continuous. Meanwhile, the input constraint is never violated due to the hard constraints in (35c).
Under the designed safety filter (35), one has
$h^{\prime}_{r}(x(t);\dot{x}(t))\geq-\alpha h_{r}(x(t))-\delta^{*}(x(t)),$ (38)
where $\dot{x}(t)=f_{0}(x(t),u^{*}(x(t)))$. When $\delta^{*}(x(t))\equiv 0$, the analysis is obvious. In fact, from (38), $h^{\prime}_{r}(x(t);\dot{x}(t))\geq-\alpha h_{r}(x(t))$, by referring to Lemma 1, $h_{r}(x(t))\geq 0$, i.e., $x(t)\in\mathcal{S},\forall t\geq 0$. Recalling (12), the volume of $\Phi(x(t))$ is always greater than a threshold $\epsilon$. ∎
Remark 4.
In Theorem 3, two conditions in (36) are attached for the sake of completeness. They could be further relaxed. i) Known from (14), the inequality in (38) can be equivalently rewritten as:
${r^{*}}^{\prime}(x(t);\dot{x}(t))\geq-\alpha{r^{*}}(x(t))+{\alpha\epsilon_{0}-\delta^{*}(x(t))}.$ (39)
This reveals that if
$\|\delta^{*}(x(t))\|_{\mathcal{L}_{\infty}}:=\sup_{t}\|\delta^{*}(x(t)\|<\alpha\epsilon_{0},$
then ${r^{*}}^{\prime}(x(t);\dot{x}(t))>-\alpha{r^{*}}(x(t))$, which means that
${r^{*}}(x(t))>0,\forall t\geq 0.$
[…]
V Numerical Simulation Example
In this section, a case study of feasible space monitoring of multiple CBFs as in Example 1, is carried out to evaluate the performance of the proposed safety filter (35).
1) Simulation Setting: Consider the scenario in Fig. 3 for a vehicle to perform a reach-avoid task under multiple obstacles and input constraints, in which feasible space monitoring of multiple CBFs is required. The vehicle is modeled as a nonholonomic dynamic unicycle:
$\dot{p}_{x}=v\cos\theta,~{}\dot{p}_{y}=v\sin\theta,~{}\dot{v}=a,~{}\dot{\theta}=\omega.$
Variables $x=[{p}_{x},{p}_{y},v,\theta]^{\top}$ are states, and $u=[a,\omega]^{\top}$ are control inputs, restricted in $\mathcal{U}=[0,2]\times[0,2]$. The reference controller in [9], with parameters $k_{p},k_{\theta},k_{v}$, is applied as the nominal control $u_{0}$ to drive the vehicle from the initial point to the goal. The barrier function candidate is designed as
$h_{i}(x)=\|[{p}_{x},{p}_{y}]^{\top}-[{o}_{xi},{o}_{yi}]^{\top}\|^{2}-R^{2},$
where $[{o}_{xi},{o}_{yi}]^{\top}$ is the center of the $i$-th obstacle and $R$ is the pre-defined safety distance. The positions of obstacles, initial and goal points can be seen from Fig. 3. The HOCBF [26] (with order $2$ and parameters $\alpha_{1}=10.0,\alpha_{2}=6.0$) is utilized to transform the position constraints $h_{i}(x)\geq 0$ to the affine constraints on input $u$, denoted as $u\in\mathcal{U}_{c}(x)$. According to Example 1, $\Phi(x)=\Psi(x)=\mathcal{U}\cap\mathcal{U}_{c}(x(t))$, the proposed safety filter (35) can be applied to solve the PVM problem here.
To show the effectiveness of (35) in improving the feasibility, it is compared with the usual safety filter that does not have the capability of feasible space monitoring:
$u^{*}(x)=$ $\operatorname*{argmin}_{u\in\Psi(x)}~{}(u-u_{0})^{\top}Q(u-u_{0}).$ (40)
In the comparison, the same nominal control $u_{0}$ is applied with $k_{p}=0.5,k_{\theta}=1$, and the same matrix $Q=[10,0;0,1]$ is applied in (35) and (40). Other parameters only appeared in (35) are set as $\epsilon_{0}=0.6$, $\alpha=15$ and $\gamma=500$.
2) Simulation Results: The numerical simulation is conducted on a laptop with an Intel Core i7-10750H CPU and 16 GB RAM. A Python simulation environment is built, similar to the one in [9]. In [9], the gradient of implicit NBF candidate $r^{*}(x)$ is approximated by some heuristic gradient value. In this paper, the in-depth analysis is performed to design an alternative one (35) without the need of the gradient of $h_{r}(x)$.
(a) The trajectories in the $XY$ plane.
(b) The time response of $r^{*}(x(t))$.
(c) The time responses of $\delta^{*}(x(t))$ and $u^{*}(x(t))$.
Figure 3: Simulation results in different scenarios: Case 1 ($k_{v}=0.5$), Case 2 ($k_{v}=1.0$), and Case 3 ($k_{v}=2.0$). Subplots (a), (b) illustrate the comparison between the proposed safety filter (35), using blue solid line, and the usual one (40) described by the green dashed line. The orange line indicates the value of user-defined threshold value $\epsilon_{0}=0.6$. Subplot (c) shows the boundedness of slack variable $\delta^{*}(x(t))$ and the continuity of inputs $u^{*}(x(t))=[a(x(t)),\omega(x(t))]^{\top}$ under the proposed safety filter (35).
To evaluate the performance sufficiently, the nominal control law $u_{0}$ with three different values of the gain $k_{v}$ are introduced in the simulation so as to produce different moving behavior, named as Case 1 ($k_{v}=0.5$), Case 2 ($k_{v}=1.0$), and Case 3 ($k_{v}=2.0$), as seen in Fig. 3.
Fig. 3(a) shows the trajectories of the vehicle in the $XY$ plane. It can be observed that under the proposed safety filter (35), the vehicle successfully avoids the obstacles and reaches the goal in all cases. Under the method being compared (40), in Case 2, the vehicle can successfully avoid the obstacles and reaches the goal, while in Case 1 and Case 3, the vehicle would get stuck in the obstacles. Fig. 3(b) illustrates the time response of $r^{*}(x(t))$, which demonstrates the range of the feasible space of these two different safety filters. From it, one can observe that $r^{*}(x(t))$ is always greater than zero and greater than $\epsilon_{0}=0.6$ in most cases, which shows that the proposed method can maintain feasibility. As a comparison, the traditional one (40) becomes infeasible in Case 1 and Case 3, and its feasible space is much smaller than the proposed one’s in Case 2. In Fig. 3(c), the time responses of slack variable $\delta^{*}(x(t))$ and inputs $u^{*}(x(t))=[a(x(t)),\omega(x(t))]^{\top}$ under the proposed safety filter (35) in the above three cases are shown. One can observe that in all cases, $\delta^{*}(x(t))$ is equal to 00 approximately and much lower than the threshold $\alpha\epsilon_{0}=9$. Besides, the inputs $u^{*}(x(t))$ are continuous in all cases. These observations confirm the reasonability of the theoretical assumptions, analysis and conclusions in Theorem 3 and Remark 4.
VI Conclusion
This paper formulates a class of control problems, called polytope volume monitoring (PVM), for which a parametric linear program-based control barrier function approach is proposed. This is achieved by establishing a bridge between the existing nonsmooth CBF and parametric optimization theory, where some useful lemmas about the continuity and directional differentiability of parametric programs are utilized to build nonsmooth CBF conditions, based on which a slacked QP-based safety filter is proposed to tackle the PVM problem step-by-step. The biggest limitation of this work is that the continuity of the proposed safety filter is not discussed formally. Future work involves designing continuous safety filters to tackle PVM problems, extending the proposed analysis and synthesis framework to other relevant problems and so on.
-A Nonsmooth Parametric Optimization
Let $\mathbb{T}\subset\mathbb{R}^{n}$ denote some open parameter set. For all $\bar{x}\in\mathbb{T}$, consider the convex parametric program with only inequalities:
$v(\bar{x}):=\min_{\bar{z}}$ $~{}\tilde{f}(\bar{z},\bar{x}),$ (A.1)
s.t. $\bar{z}\in{F}(\bar{x}):=\left\{\bar{z}\in\mathbb{R}^{n_{z}}\mid\tilde{g}_{j}(\bar{z},\bar{x})\leq 0,j\in\mathcal{J}\right\},$
with $\mathcal{J}:=\{1,2,\ldots,N\}$. For convenience, let the set of (global) minimizers according to (A.1) be denoted as
$S(\bar{x}):=\left\{\bar{z}\in F(\bar{x})\mid\tilde{f}(\bar{z},\bar{x})=v(\bar{x})\right\}.$ (A.2)
For $\bar{x}\in\mathbb{T}$ and feasible $\bar{z}\in F(\bar{x})$, the active index set is defined as $\mathcal{J}_{0}(\bar{z},\bar{x}):=\left\{j\in\mathcal{J}\mid\tilde{g}_{j}(\bar{z},\bar{x})=0\right\}$. Throughout this section, $\tilde{f},\tilde{g}_{j},j\in\mathcal{J}$, are continuously differentiable on $\mathbb{R}^{n_{z}}\times\mathbb{T}$, notation $\tilde{f},\tilde{g}_{j}\in\mathcal{C}^{1}$. The following Constraint Qualification conditions can be found in [18, (1.3) and (1.4)] and [27, (11) and (12)].
Definition 3.
There are two constraint qualification conditions, respectively defined as:
1) The Mangasarian Fromovitz Constraint Qualification (MFCQ) is said to hold at $(\bar{z},\bar{x})$ for $\bar{z}\in F(\bar{x})$ if there exists $\xi$ such that $\partial_{z}\tilde{g}_{j}(\bar{z},\bar{x})\xi<0,j\in\mathcal{J}_{0}(\bar{z},\bar{x})$.
2) Linear Independence Constraint Qualification (LICQ) holds at $(\bar{z},\bar{x})$ if vectors $\partial_{z}\tilde{g}_{j}(\bar{z},\bar{x}),j\in\mathcal{J}_{0}(\bar{z},\bar{x})$ are linearly independent.
The continuity and directional differentiability properties of the value function of parametric programs with only inequalities have been studied in [18].
Lemma 2.
[18, Lemma 6.2] Consider the parametric convex program in (A.1) and assume there exists a neighborhood $B_{\varepsilon_{0}}(\bar{x}),\varepsilon_{0}>0$, such that $\cup_{x\in B_{\varepsilon_{0}}(\bar{x})}F(x)\subset C_{0}$, with $C_{0}$ compact. If the MFCQ condition holds at $(\bar{z},\bar{x})$ for any $\bar{z}\in S(\bar{x})$, then the value function $v(\cdot)$ is locally Lipschitz continuous at $\bar{x}$.
Define the (local) Lagrangian function near $(\bar{z},\bar{x})$ as:
$L(\mu,z,x)=\tilde{f}(z,x)+\sum_{j\in\mathcal{J}_{0}(\bar{z},\bar{x})}\mu_{j}\tilde{g}_{j}(z,x),$ (A.3)
where ${\mu}:=\left({\mu}_{j},j\in\mathcal{J}_{0}(\bar{z},\bar{x})\right)\in\mathbb{R}^{\left|\mathcal{J}_{0}(\bar{z},\bar{x})\right|}$, collecting all multiplers $\mu_{j}$ corresponding to active constraints in $\mathcal{J}_{0}(\bar{z},\bar{x})$. Then the partial derivatives of $L(\cdot)$ are respectively given by
$\partial_{x}L(\mu,z,x):=$ $\partial_{x}\tilde{f}(z,x)+\sum_{j\in\mathcal{J}_{0}(\bar{z},\bar{x})}\mu_{j}\partial_{x}\tilde{g}_{j}(z,x),$
$\partial_{z}L(\mu,z,x):=$ $\partial_{z}\tilde{f}(z,x)+\sum_{j\in\mathcal{J}_{0}(\bar{z},\bar{x})}\mu_{j}\partial_{z}\tilde{g}_{j}(z,x).$
Lemma 3.
[18, Theorem 7.3] Consider the parametric convex program in (A.1) and let $F(\bar{x})$ be compact. If the MFCQ condition holds at $(\bar{z},\bar{x})$ satisfying $\bar{z}\in F(\bar{x})$, then $v(\cdot)$ is directional differentiable at $\bar{x}$ in any direction $d,\|d\|=1$ with
$v^{\prime}(\bar{x};d)=\min_{\bar{z}\in S(\bar{x})}\max_{\mu\in M(\bar{z},\bar{x})}\partial_{x}L(\mu,\bar{z},\bar{x})d,$ (A.5)
where the minimizer set $S(\cdot)$ is defined in (A.2), and
$M(\bar{z},\bar{x})$ :=$:=\left\{\mu\mid\partial_{z}L(\mu,\bar{z},\bar{x})=0^{\top},\mu\geq 0\right\}.$ (A.6)
If LICQ holds, then $M(\bar{z},\bar{x})$ will be a singleton, i.e., the multiplier ${\mu}$ will be uniquely determined.
The min-max expression of directional derivative in (A.5) has also been pointed out by other literature [28, Theorem 2.4], [29, Theorem 4.24] yet under different conditions.
-B Proofs
Proposition 1 can be proved by some standard results in LP theory. Proposition 2 is the corollary of Lemma 3 with the boundness of set $M(z,x)$ to be proven.
Proof of Proposition 1: i) First, it is shown that $F_{P}(x)\neq\emptyset$ for all $x$. Recalling
$\Phi(x)=\{c\mid a_{j}^{\top}(x)c\leq b_{j}(x),~{}\forall j\in\mathcal{J}\},$
is nonempty, compact and convex, one can find a sufficiently small scalar $\varepsilon_{0}(x)>0$ and a point $c_{0}(x)$, such that
$a_{j}^{\top}(x)c_{0}(x)+\varepsilon_{0}(x)\leq b_{j}(x),~{}\forall j=1,2,\cdots,N.$
For every $j$, since $\|a_{j}(x)\|\neq 0$, there exists $c_{2}(x)\neq 0$, s.t.
$\varepsilon_{0}=a_{j}^{\top}c_{2}=\|a_{j}\|\underbrace{\|c_{2}\|\cos\langle a_{j},c_{2}\rangle}_{=:r_{j}(x)}>0.$
Defining $r_{0}:=\min_{j\in\mathcal{J}}~{}r_{j}$, one has $r_{0}>0$ and $\|a_{j}\|r_{0}\leq\varepsilon_{0}$, which leads to
$a_{j}^{\top}(x)c_{0}(x)+\|a_{j}(x)\|r_{0}(x)\leq b_{j}(x),~{}\forall j=1,2,\cdots,N,$
i.e., $F_{P}(x)\neq\emptyset$. Besides, $\Phi(x)$ is uniformly bounded for all $x$ as in Assumption 1, thus $F_{P}(x)$ is also uniformly bounded.
ii) Since the feasible set $F_{P}(x)$ is uniformly bounded, $r^{*}(x)$ and $h_{r}(x)$ are also uniformly bounded. Due to the fact that (14) is a LP problem, the optimization value is obtained at a point $z$ only if $z$ is on the boundary $\partial F_{P}(x)$, hence $S(x)\subset\partial F_{P}(x)$ is also uniformly bounded.
[…]
Proof of Proposition 2: According to (A.3), the local Lagrangian function near $\bar{x},\bar{z}$ is defined as:
$L(\mu,z,x)=q^{\top}z+\mu^{\top}(\tilde{\mathbf{A}}_{\mathcal{J}_{0}(\bar{z},\bar{x})}(x)z-\tilde{\mathbf{b}}_{\mathcal{J}_{0}(\bar{z},\bar{x})}(x)),$
where $\mu\in\mathbb{R}^{|\mathcal{J}_{0}({z},x)|}$, and $\tilde{\mathbf{A}}_{\mathcal{J}_{0}},\tilde{\mathbf{b}}_{\mathcal{J}_{0}}$ are defined in (27). Then, the partial derivative of $L(\mu,z,x)$ w.r.t. vectors $z$ and $x$ are respectively given by:
$\partial_{z}L(\mu,z,x)$ $=q^{\top}+\mu^{\top}\tilde{\mathbf{A}}_{\mathcal{J}_{0}(\bar{z},\bar{x})}(x),$
$\partial_{x}L(\mu,z,x)$ $=\mu^{\top}\left(\frac{\partial}{\partial x}\left(\tilde{\mathbf{A}}_{\mathcal{J}_{0}(\bar{z},\bar{x})}(x)z\right)-\frac{\partial}{\partial x}\tilde{\mathbf{b}}_{\mathcal{J}_{0}(\bar{z},\bar{x})}(x)\right)$
$=\sum_{j\in\mathcal{J}_{0}(\bar{z},\bar{x})}\mu_{j}\left(z^{\top}\frac{\partial\tilde{{a}}_{j}}{\partial x}(x)-\frac{\partial{{b}}_{j}}{\partial x}(x)\right),$
and $M(z,x):=\left\{\mu\mid\partial_{z}L(\mu,z,x)=0^{\top},~{}\mu\geq 0\right\}.$ In the above, the partial derivative of $L(\mu,z,x)$ w.r.t. variable $z,x$ at the point $\bar{z},\bar{x}$ are calculated. By further choosing $\bar{z}=z,\bar{x}=x$, the expressions in (29) and (30) can be obtained.
Since MFCQ condition holds and $F_{P}(x)$ is compact as proved in Proposition 1, then the conditions of Lemma 3 hold, from which one has
$h^{\prime}_{r}(x;\dot{x})$ $=-v^{\prime}_{r}(x;\dot{x})$
$=-\min_{z\in S(x)}\max_{\mu\in M(z,x)}\partial_{x}L(\mu,z,x)\dot{x}$
$=\max_{z\in S(x)}\min_{\mu\in M(z,x)}-\partial_{x}L(\mu,z,x)\dot{x}.$
Hence, the expression in (28) is obtained.
The set $M(z,x)$ in (30) can be further rewritten as:
$M(z,x)$ =$=\left\{\mu\mid\sum_{j\in\mathcal{J}_{0}({z},{x})}\tilde{a}_{j}(x)\mu_{j}=-q,~{}\mu\geq 0\right\}$
$=M_{1}(z,x)\cap M_{2}(z,x),$
where
$M_{1}(z,x)$ :=$:=\left\{\mu\mid\sum_{j\in\mathcal{J}_{0}({z},{x})}a_{j}(x)\mu_{j}=0\right\},$
$M_{2}(z,x)$ :=$:=\left\{\mu\mid\sum_{j\in\mathcal{J}_{0}({z},{x})}\|a_{j}(x)\|\mu_{j}=1,~{}\mu\geq 0\right\}.$
[…]

### v3

@@
 Recent studies have explored the idea of applying Chebyshev ball method in optimization theory to solve the case study of PVM; however, the underlying difficulties caused by nonsmoothness have not been addressed.
-This paper continues the study on this topic, where our main contribution is to establish the relationship between nonsmooth CBF and parametric optimization theory through directional derivatives for the first time, so as to solve PVM problems more conveniently.
+This paper continues the study on this topic, where our main contribution is to establish the relationship between nonsmooth CBF and parametric optimization theory through directional derivatives for the first time, to solve PVM problems more conveniently.
 In detail, inspired by Chebyshev ball approach, a parametric linear program (PLP) based nonsmooth barrier function candidate is established for PVM, and then, sufficient conditions for it to be a nonsmooth CBF are proposed, based on which a quadratic program (QP) based safety filter with guaranteed feasibility is proposed to address PVM problems.
@@
 Meanwhile, a simpler way of addressing the feasibility problem appears in [9].
-In detail, “Feasible CBF Space” is defined to describe the state-dependant polytope (in input space) formulated by CBF conditions and input constraints.
-Then, Feasible Space Monitoring (FSM) problem is established to monitor the feasibility of the resulting QP-based control laws, for which Chebyshev ball and inscribing ellipsoid methods are proposed to build parametric linear program (PLP) based CBFs.
+In detail, “Feasible CBF Space” is defined to describe the state-dependant polytope formulated by CBF conditions and input constraints.
+Then, “Feasible Space Monitoring” (FSM) problem is established to monitor the feasibility of the resulting QP-based control laws, for which Chebyshev ball and inscribing ellipsoid methods are proposed to build parametric convex program based CBFs.
 For engineering implementations, the differentiable optimization toolbox is applied to approximate the gradient of the resulting (nonsmooth implicit) barrier function candidate by some heuristic gradient values.
@@
 Such observation motivates this paper.
-Meanwhile, researchers have been gradually interested in building CBFs via parametric optimization-defined metrics for related problems, such as network connectivity maintenance [10] and polytopic collision avoidance (PCA) [11, 12, 13, 14, 15, 16].
+Meanwhile, researchers have been gradually interested in building CBFs via parametric optimization-defined metrics for other problems, such as network connectivity maintenance [10] and polytopic collision avoidance (PCA) [11, 12, 13, 14, 15, 16].
 However, it is difficult to extend these results to solve the above open problem in FSM.
@@
 First, a broader class of control problems is formulated, called Polytope Volume Monitoring (PVM).
-By identifying this kind of problems, FSM problem [9] and Polytopic Intersectability Maintenance (PIM), i.e., the opposite problem of PCA [11, 12, 13, 14, 15], can be regarded as different special examples of PVM, as seen in Fig.
-1.
+By identifying this kind of problems, FSM problem [9] and Polytopic Intersectability Maintenance, i.e., the opposite problem of PCA [11, 12, 13, 14, 15], can be regarded as different special examples of PVM.
 This generalization enhances the applicability of the proposed method to a wider range of problems.
-Second, to address difficulties in analysis and synthesis caused by inherent nonsmoothness, this paper attempts to establish a bridge between the directional derivative defined nonsmooth CBF [17] and parametric optimization theory [18].
+Second, to address difficulties in analysis and synthesis caused by inherent nonsmoothness, this paper attempts to establish a bridge between the directional-derivative-defined nonsmooth CBF [17] and parametric optimization theory [18].
 In detail, our main contributions can be summarized as follows:
@@
   • 2)
-    Within this unified framework, a parametric LP based nonsmooth barrier function candidate is established for PVM, based on which sufficient conditions of being a nonsmooth CBF are proposed.
-Then a slacked QP-based safety filter with guaranteed feasibility is proposed to solve the PVM problem.
+    Within this unified framework, a parametric linear program based nonsmooth barrier function candidate is established for PVM, based on which sufficient conditions of being a nonsmooth CBF are proposed.
+Then a a feasibility-guaranteed QP-based safety filter is proposed to solve the PVM problem.
 The above nonsmooth analysis and considerations are not addressed in [9].
@@
 Section IV reports the main results.
-Numerical simulation is performed in Section V.
-Appendix includes the referred nonsmooth parametric optimization theory and proofs.
-Notations: Throughout this paper, $\operatorname{col}(x_{1},x_{2}):=[x_{1}^{\top},x_{2}^{\top}]^{\top}$ for any two column vectors $x_{1},x_{2}$.
+The referred nonsmooth parametric optimization theory is included in Appendix.
+Notations: Throughout this paper, $\operatorname{col}(x_{1},x_{2})\mathrel{\mathop{\ordinarycolon}}=[x_{1}^{\top},x_{2}^{\top}]^{\top}$ for any two column vectors $x_{1},x_{2}$.
 For any vector $x$, $\|x\|$ denotes its Euclidean norm.
@@
 The interior and boundary of a set $\mathcal{C}$ are indicated by $\operatorname{Int}(\mathcal{C})$ and $\partial\mathcal{C}$.
-A continuous function $\alpha:\mathbb{R}\rightarrow\mathbb{R}$ is called in extended class-$\mathcal{K}$ if it is strictly increasing and with $\alpha(0)=0$ [1].
-For any function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$, if exists, its generalized gradient at $x$ is denoted by $\partial f(x)$, and $f^{\prime}(x;d)$ denotes its directional derivative at $x$ along direction $d$.
+A continuous function $\alpha\mathrel{\mathop{\ordinarycolon}}\mathbb{R}\rightarrow\mathbb{R}$ is called in extended class-$\mathcal{K}$ if it is strictly increasing and with $\alpha(0)=0$ [1].
+For any function $f\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{n}\rightarrow\mathbb{R}$, if exists, its generalized gradient at $x$ is denoted by $\partial f(x)$, and $f^{\prime}(x;d)$ denotes its directional derivative at $x$ along direction $d$.
 For brevity, its partial derivative $\frac{\partial f}{\partial x}(x)$, expressed as a row vector, is denoted by $\partial_{x}f$ sometimes.
@@
 Consider a control-affine nonliear system
-$\dot{x}(t)=f_{0}(x(t),u(t)):=f(x(t))+g(x(t))u(t),$ (1)
+$\dot{x}(t)=f_{0}(x(t),u(t))\mathrel{\mathop{\ordinarycolon}}=f(x(t))+g(x(t))u(t),$ (1)
 where $x\in\mathbb{R}^{n}$ is state, $u\in\mathcal{D}_{u}\subset\mathbb{R}^{m}$ is control input, and $f,g$ are Lipschitz continuous.
@@
 Definition 1.
-Given a nonempty closed set $\mathcal{C}\subset\mathbb{R}^{n}$ with no isolated point, a locally Lipschitz continuous function $h$ : $\mathbb{R}^{n}\rightarrow\mathbb{R}$ is called a NBF candidate on set $\mathcal{C}$ if
+Given a nonempty closed set $\mathcal{C}\subset\mathbb{R}^{n}$ with no isolated point, a locally Lipschitz continuous function $h(\cdot)\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{n}\rightarrow\mathbb{R}$ is called a NBF candidate on set $\mathcal{C}$ if
 $h(x)$ $=0,~\forall x\in\partial\mathcal{C};\quad h(x)>0,~\forall x\in\operatorname{Int}(\mathcal{C});$ (2)
 $h(x)$ $<0,~\forall x\in\mathbb{R}^{n}/\mathcal{C}.$
-Function $h$ is called a differentiable CBF to (1) if BF candidate $h$ is further differentiable and there exists an enxtended $\mathcal{K}$-class function $\alpha$ such that for all $x$ it holds $\sup_{u\in\mathcal{D}_{u}}\left\{\frac{\partial h}{\partial x}(x)f_{0}(x,u)\right\}\geq-\alpha(h(x)).$ For the demands of this paper, we only require from now on that $h$ is Lipschitz continuous and the (one-sided) directional derivative of $h(x)$ at point $x$ along the direction $\dot{x}$ exists, i.e., the following limit is finite and abbreviated as
-$h^{\prime}(x(t);\dot{x}(t))$ $:=\lim_{\sigma\rightarrow 0^{+}}\frac{h(x(t)+\sigma\dot{x}(t))-h(x(t))}{\sigma}.$ (3)
+For the demands of this paper, we only require from now on that $h$ is locally Lipschitz continuous and the (one-sided) directional derivative of $h(x)$ at point $x$ along any direction $d$ exists, i.e., the following limit is finite and abbreviated as
+$h^{\prime}(x;d)\mathrel{\mathop{\ordinarycolon}}=\lim_{\sigma\rightarrow 0^{+}}\frac{h(x+\sigma d)-h(x)}{\sigma}.$
 By doing this, one can relax the differentiability of $h$ and define a broader class of CBFs as follows.
 Definition 2.
-A locally Lipschitz continuous function $h:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is a nonsmooth control barrier function (NCBF) for system (1) if $h$ is a NBF candidate on the closed set $\mathcal{C}\subset\mathbb{R}^{n}$, and there exists an extended class-$\mathcal{K}$ function $\alpha:\mathbb{R}\rightarrow\mathbb{R}$, an open set $\mathcal{D}$ that satisfies $\mathcal{C}\subset\mathcal{D}\subseteq\mathbb{R}^{n}$, such that:
+A locally Lipschitz continuous function $h\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{n}\rightarrow\mathbb{R}$ is a nonsmooth control barrier function (NCBF) for system (1) if $h$ is a NBF candidate on the closed set $\mathcal{C}\subset\mathbb{R}^{n}$, and there exists an extended class-$\mathcal{K}$ function $\alpha\mathrel{\mathop{\ordinarycolon}}\mathbb{R}\rightarrow\mathbb{R}$, an open set $\mathcal{D}$ that satisfies $\mathcal{C}\subset\mathcal{D}\subseteq\mathbb{R}^{n}$, such that:
 $\sup_{u\in\mathcal{D}_{u}}\left\{h^{\prime}\left(x;f_{0}(x,u)\right)\right\}\geq-\alpha(h(x)),\quad\forall x\in\mathcal{D}.$ (3)
-In [17], the same limitation value as $h^{\prime}(x;f_{0}(x,u))$ is used to define NCBF, yet without introducing directional derivative formally.
+In [17], the same limitation value as $h^{\prime}(x;f_{0}(x,u))$ is used to define NCBF, yet without introducing the concept of directional derivative.
 In this paper, the directional derivative defined NCBF (3) is introduced to build a bridge between the NCBF theory and the parametric optimization theory.
 Lemma 1.
-[17, Theorem 1]  If $h$ is a NCBF for system (1), then the closed set $\mathcal{C}$ is forward control-invariant, that is, there exists $u(t)\in\mathcal{D}_{u}$, such that $x_{0}\in\mathcal{C}$ implies that $x(t)\in\mathcal{C},~{}\forall t\in\left[0,T\right]$, for every solution.
+[17, Theorem 1]  If $h$ is a NCBF for system (1), then the closed set $\mathcal{C}$ is forward control-invariant, that is, there exists a measurable and locally bounded function $u(t)\in\mathcal{D}_{u}$, such that $x_{0}\in\mathcal{C}$ implies that $x(t)\in\mathcal{C},~\forall t\in\left[0,T\right]$ for every solution.
 When $h(x)$ is differentiable, it holds that $h^{\prime}(x;f_{0}(x,u))=\frac{\partial h}{\partial x}(x)\left(f(x)+g(x)u\right)$, which is a linear function of input $u$ at every $x$, and the following is a QP-defined control law:
@@
 where $u_{0}$ is any given continuous nominal control law.
-When $h(x)$ is no longer differentiable but directional derivative $h^{\prime}(x;f_{0}(x,u))$ exists, the inequality (5b) may no longer be a linear constraint of input $u$.
+When $h(x)$ is no longer differentiable but the directional derivative $h^{\prime}(x;f_{0}(x,u))$ exists, the inequality (4b) may no longer be a linear constraint of input $u$.
 Remark 1.
@@
 […]
-III Problem Formulation and Analysis
+(a) Polytope volume monitoring problem and examples.
+(b) The structure of the proposed safety filter.
+Figure 1: A summary of the polytope volume monitoring problem and the parametric linear program based control barrier function.
+III Problem Formulation
 III-A Problem Statement
@@
 $\dot{x}$ $=f(x)+g(x)u,$ (6a)
-$y$ $=\Phi(x):=\left\{c\in\mathbb{R}^{l}\mid a_{j}(x)^{\top}c\leq b_{j}(x),j\in\mathcal{J}\right\},$ (7b)
-$u$ $\in\Psi(x):=\left\{u\in\mathbb{R}^{m}\mid\mathfrak{a}_{\mathfrak{i}}(x)^{\top}u\leq\mathfrak{b}_{\mathfrak{i}}(x),\mathfrak{i}\in\mathfrak{J}\right\},$ (7c)
-where $\mathcal{J}:=\{1,2,...,N\}$ and $\mathfrak{J}:=\{1,2,...,\mathfrak{N}\}$.
-For convenience, the functions appearing in definitions of $\Phi(x)$ and $\Psi(x)$ are assumed to be continuously differentiable, and the following assumptions are imposed to develop an implementable controller.
+$y$ $=\Phi(x)\mathrel{\mathop{\ordinarycolon}}=\left\{c\in\mathbb{R}^{l}\mid a_{j}^{\top}(x)c\leq b_{j}(x),j\in\mathcal{J}\right\},$ (6b)
+$u$ $\in\Psi(x)\mathrel{\mathop{\ordinarycolon}}=\left\{u\in\mathbb{R}^{m}\mid\mathfrak{a}_{\mathfrak{i}}^{\top}(x)u\leq\mathfrak{b}_{\mathfrak{i}}(x),\mathfrak{i}\in\mathfrak{J}\right\},$ (6c)
+where $\mathcal{J}\mathrel{\mathop{\ordinarycolon}}=\{1,2,...,N\}$ and $\mathfrak{J}\mathrel{\mathop{\ordinarycolon}}=\{1,2,...,\mathfrak{N}\}$.
+For convenience, the functions appearing in definitions of $\Phi(x)$ and $\Psi(x)$ are assumed to be continuously differentiable, and without loss of generality, the following assumptions are imposed to develop an implementable controller.
 Assumption 1.
-The state-dependent polytopic sets $\Phi(x)$, $\Psi(x)$ are both nonempty and uniformly bounded for all $x$, i.e., there exist compact sets $\mathcal{D}_{y}\subset\mathbb{R}^{l}$ and $\mathcal{D}_{u}\subset\mathbb{R}^{m}$, such that $\Phi(x)\subseteq\mathcal{D}_{y}$ and $\Psi(x)\subseteq\mathcal{D}_{u}$ respectively.
-Additionally, vectors $a_{j}(x),j\in\mathcal{J}$ are nonzero at any $x$.
-Without loss of generality, Assumption 1 is made for some technical considerations.
-Especially, the non-emptiness of $\Phi(x)$ is essential to guarantee that some metric is well-defined in the following sections.
+The polytopic input constraint set $\Psi(x)$ is always nonempty and the vectors $a_{j}(x),j\in\mathcal{J}$ are always nonzero.
+The state-dependent polytopic set $\Phi(x)$ is uniformly bounded, i.e., there exists a compact set $\mathcal{D}_{y}\subset\mathbb{R}^{l}$, such that $\Phi(x)\subseteq\mathcal{D}_{y},\forall x\in\mathbb{R}^{n}$.
+It is worth noting that any emptyset is naturally bounded, thus it is not required in Assumption 1 that $\Phi(x)$ is always nonempty.
+Instead, we allow there exist some points $x$ at which $\Phi(x)=\emptyset$.
 In fact, this paper will attempt to synthesize a controller to prevent the interior of $\Phi(x(t))$ from vanishing, as formally stated below.
-Figure 1: Illustration of the PVM problem and its examples. (a) Problem statement and preliminary analysis. (b) Example 1 in input space. (c) Example 2 in state space.
 Polytope Volume Monitoring Problem: Consider the system (6) with a polytopic set-valued output $\Phi(x)$ and a polytopic input constraint set $\Psi(x)$.
-The goal is to design a control law for $u\in\Psi(x)$ to prevent the volume of the polytope $\Phi(x(t))$ from being less than a pre-defined positive value, i.e.,
-$\mathcal{V}(x(t)):=\int_{\Phi(x(t))}\mathrm{d}s\geq\epsilon,~{}\forall t\geq 0,$ (8)
-where $\epsilon>0$ is any user-defined threshold value.
-There are some typical situations where the volume of state-dependant polytope $\Phi(x(t))$ might be zero if the input $u$ is not designed properly.
+The goal is to design a control law for $u$ to prevent the volume of $\Phi(x(t))$ from decreasing to zero, i.e., to ensure the nonemptness of the interior of $\Phi(x(t))$:
+$x(t)\in\mathcal{A}\mathrel{\mathop{\ordinarycolon}}=\{x\mid\operatorname{Int}(\Phi(x))\neq\emptyset\},\quad\forall t\geq 0.$ (7)
+There are some typical situations where the state-dependant polytope $\Phi(x(t))$ might be empty if the input $u$ is designed improperly.
 Example 1 (Feasible Space Monitoring [9]).
 In Fig.
-1(b), given the polytopic input constraint set $\mathcal{U}$, by constructing the set composed of multiple differentiable CBFs $h_{i},i\in\mathcal{I}$: $\mathcal{U}_{c}(x):=\{u\mid\frac{\partial h_{i}}{\partial x}(x)\left(f(x)+g(x)u\right)\geq-\alpha_{i}(h_{i}(x)),i\in\mathcal{I}\}$, the FSM problem aims to design $u$ such that
-$\mathcal{U}\cap\mathcal{U}_{c}(x(t))\neq\emptyset,~{}\forall t.$ (9)
-By choosing $\Phi(x)=\Psi(x)=\mathcal{U}\cap\mathcal{U}_{c}(x)$, feasible space monitoring of any QP-based controllers can be regarded as case studies of the formulated PVM problem.
+1 (a), given the polytopic input constraint set $\mathcal{U}$, by constructing the set composed of multiple differentiable CBFs $h_{i},i\in\mathcal{I}$: $\mathcal{U}_{c}(x)\mathrel{\mathop{\ordinarycolon}}=\{u\mid\frac{\partial h_{i}}{\partial x}(x)\left(f(x)+g(x)u\right)\geq-\alpha_{i}(h_{i}(x)),i\in\mathcal{I}\}$, the FSM problem aims to design $u$ such that
+$\mathcal{U}\cap\mathcal{U}_{c}(x(t))\neq\emptyset,~\forall t\geq 0.$ (8)
+By choosing $\Phi(x)=\Psi(x)=\mathcal{U}\cap\mathcal{U}_{c}(x)$, FSM of any QP-based controllers can be regarded as the case studies of the formulated PVM problem.
 In addition to the above examples in input space, PVM problems also appear in state space.
@@
 Consider two standard nonholonomic vehicles $i,j$, for each $\mathfrak{i}\in\{i,j\}$, $x^{\mathfrak{i}}=[{p}_{x}^{\mathfrak{i}},{p}_{y}^{\mathfrak{i}},{\theta}^{\mathfrak{i}}]^{\top}$ and $u^{\mathfrak{i}}=[v^{\mathfrak{i}},\omega^{\mathfrak{i}}]^{\top}\in\mathcal{U}^{\mathfrak{i}}$ are state and input vectors respectively, where $\mathcal{U}^{\mathfrak{i}}$ denotes a hyperrectangle.
-Let sets $\mathcal{P}^{i}(x^{i})$ and $\mathcal{P}^{j}(x^{j})$ represent their physical domain or sensor ranges respectively, as shown in Fig.
-1(c).
-The goal of polytopic intersectability maintenance is to maintain the intersection of two polytopes nonempty:
-$\mathcal{P}^{i}(x^{i}(t))\cap\mathcal{P}^{j}(x^{j}(t))\neq\emptyset,~{}\forall t,$ (10)
+Let $\mathcal{P}^{i}(x^{i}),\mathcal{P}^{j}(x^{j})$ denote their physical domain or sensor ranges, as in Fig.
+1 (a).
+The goal of PIM is to maintain the intersection of two polytopes nonempty:
+$\mathcal{P}^{i}(x^{i}(t))\cap\mathcal{P}^{j}(x^{j}(t))\neq\emptyset,~\forall t\geq 0,$ (9)
 which is opposite to the goals of polytopic collision avoidance [11, 12, 13, 14, 15, 16].
@@
 III-B Preliminary Analysis
-The center $c\in\mathbb{R}^{l}$ and radius $r\geq 0$ of the largest sphere contained inside the polytope $\Phi(x)$ defined in (7b), called Chebyshev ball [21, Chapter 8], can be computed by the following Linear Program at every $x$:
-$r^{*}(x):=$ $\max_{c,r}~{}r=-\min_{c,r}~{}-r,$ (11a)
-s.t. $a_{j}^{\top}(x)c+\|a_{j}(x)\|r\leq b_{j}(x),~{}\forall j\in\mathcal{J},$ (11b)
-then from the geometric meaning in Fig. 1(a), it holds that
-$\mathcal{V}(x(t))\geq\pi(r^{*}(x(t)))^{2}\geq\epsilon,$ (12)
-from which one knows the control goal in (8) would hold if
-$r^{*}(x(t))\geq\epsilon_{0}:=\sqrt{\frac{\epsilon}{\pi}},~{}\forall t\geq 0.$ (13)
+The radius $r\geq 0$ of the largest sphere contained inside the polytope $\Phi(x)$ defined in (6b), called Chebyshev ball [21, Chapter 8], can be computed by the following linear program (LP) at every $x$:
+$r^{*}(x)\mathrel{\mathop{\ordinarycolon}}$= $\max_{c\in\mathbb{R}^{l},r\geq 0}~r=-\min_{c\in\mathbb{R}^{l},r\geq 0}~-r,$ (10a)
+s.t. $a_{j}^{\top}(x)c+\|a_{j}(x)\|r\leq b_{j}(x),~\forall j\in\mathcal{J}.$ (10b)
+Given any user-defined threshold value $\epsilon_{0}>0$, as illustrated in Fig. 1 (a), one can know that the goal in (7) holds if
+$r^{*}(x(t))\geq\epsilon_{0},~\forall t\geq 0.$ (11)
 Remark 2.
@@
 IV Main results
-To solve the aforementioned open problem, the nonsmooth parametric optimization theory, as shown in Appendix -A, is introduced to develop a general approach for synthesizing a NCBF-based safety filter to solve PVM problems for the first time.
+To solve the aforementioned open problem, the nonsmooth parametric optimization theory [18], as shown in Appendix A-A, is introduced to develop a general approach for synthesizing a NCBF-based safety filter to solve PVM problems for the first time.
 Without calculating the generalized gradient $\partial r^{*}(x)$, the NCBF condition is established based on its directional derivative for convenience.
 IV-A Establishing a Parameteric LP-defined NBF Candidate
-Inspired by (13), one can define an implicit function $h_{r}(x)$ via the following parametric linear program:
-$h_{r}(x):=$ $r^{*}(x)-\epsilon_{0}=-\epsilon_{0}-\min_{z\in F_{P}(x)}~{}q^{\top}z,$ (14)
-where $q:=\operatorname{col}{(0_{l\times 1},-1)}$ and $z:=\operatorname{col}{(c,r)}\in\mathbb{R}^{l+1}$. The feasible set $F_{P}(x)$ is defined as
-$F_{P}(x):=\left\{z\mid\tilde{\mathbf{A}}(x)z\leq\tilde{\mathbf{b}}(x)\right\},$ (15)
-where $\tilde{\mathbf{A}}(x)\in\mathbb{R}^{N\times(l+1)}$ and $\tilde{\mathbf{b}}(x)\in\mathbb{R}^{N}$ are
-$\tilde{\mathbf{A}}(x):=\left[\begin{array}[]{cc}\tilde{a}_{1}(x)^{\top}\\ \tilde{a}_{2}(x)^{\top}\\ \cdots\cdots\\ \tilde{a}_{N}(x)^{\top}\end{array}\right],\quad\tilde{\mathbf{b}}(x):=\left[\begin{array}[]{c}b_{1}(x)\\ b_{2}(x)\\ \cdots\cdots\\ b_{N}(x)\end{array}\right],$ (24)
-with $\tilde{a}_{j}(x):=\operatorname{col}(a_{j}(x),\|a_{j}(x)\|)$ for every $j=1,2,...,N$.
-Next, preliminary properties of the PLP in (14) are shown first, where Mangasarian Fromovitz Constraint Qualification (MFCQ), as in Definition 3, is a key condition in studying the continuity and directional differentiability of $r^{*}(x)$.
+Inspired by (10) and (11), one can define an implicit function $h_{r}(x)$ via the following parametric linear program:
+$h_{r}(x)\mathrel{\mathop{\ordinarycolon}}$= $r^{*}(x)-\epsilon_{0}=-\epsilon_{0}-\!\min_{z\in F_{P}(x)}q^{\top}z,\forall x\in\mathcal{A},$ (12)
+where $q\mathrel{\mathop{\ordinarycolon}}=\operatorname{col}{(0_{l\times 1},-1)}$ and $z\mathrel{\mathop{\ordinarycolon}}=\operatorname{col}{(c,r)}\in\mathbb{R}^{l+1}$.
+The feasible set $F_{P}(x)$ is defined as
+$F_{P}(x)\mathrel{\mathop{\ordinarycolon}}=\left\{z\mid\tilde{\mathbf{A}}(x)z\leq\tilde{\mathbf{b}}(x)\right\},$ (13)
+where $\tilde{\mathbf{A}}(x)\in\mathbb{R}^{(N+1)\times(l+1)}$ and $\tilde{\mathbf{b}}(x)\in\mathbb{R}^{N+1}$ are
+$\tilde{\mathbf{A}}(x)\mathrel{\mathop{\ordinarycolon}}=\left[\!\!\!\begin{array}[]{cc}0_{1\times l},&-1\\ a_{1}^{\top}(x),&\|a_{1}(x)\|\\ a_{2}^{\top}(x),&\|a_{2}(x)\|\\ \cdots&\cdots\\ a_{N}^{\top}(x),&\|a_{N}(x)\|\end{array}\!\!\!\right],~\tilde{\mathbf{b}}(x)\mathrel{\mathop{\ordinarycolon}}=\left[\!\!\!\begin{array}[]{c}0\\ b_{1}(x)\\ b_{2}(x)\\ \cdots\\ b_{N}(x)\end{array}\!\!\!\right].$ (24)
+For brevity, let $\tilde{a}_{j}^{\top}(x)$ and $\tilde{b}_{j}(x)$ denote the $j$-th row of $\tilde{\mathbf{A}}(x)$ and $\tilde{\mathbf{b}}(x)$ respectively, i.e.,
+$\tilde{a}_{j}(x)\mathrel{\mathop{\ordinarycolon}}=\tilde{\mathbf{A}}^{\top}(x)e_{j},~\tilde{b}_{j}(x)\mathrel{\mathop{\ordinarycolon}}=\tilde{\mathbf{b}}^{\top}(x)e_{j},\forall j\in[N+1],$
+where $e_{j}$ is the $j$-th column of the $(N+1)$-dimensional identity matrix and $N+1]\mathrel{\mathop{\ordinarycolon}}=\{1,2,\cdots,N+1\}$.
+Next, the basic properties of the parametric linear program in (12) are shown first, where Mangasarian Fromovitz Constraint Qualification (MFCQ), as in Definition A.1, is a key condition in studying the continuity and directional differentiability of $r^{*}(x)$.
 Proposition 1.
-For the value function $h_{r}(x)$ of the PLP in (14), the following conclusions hold:
+Consider the set $\mathcal{A}$ defined in (7) and the function $h_{r}(x)$ in (12). The following conclusions hold:
   • i)
-    The feasible set $F_{P}(x)$ is nonempty for every $x$ and uniformly bounded for all $x$.
+    The interior of $F_{P}(x)$ is nonempty for all $x\in\mathcal{A}$.
   • ii)
-    For every $x$, $h_{r}(x)$ is finite, and the solution map
-    $S(x):=$ $\operatorname*{argmin}_{z\in F_{P}(x)}~{}q^{\top}z,$ (25)
+    Furthermore, the set $F_{P}(x)$ and the function $h_{r}(x)$ are uniformly bounded for all $x\in\mathcal{A}$. The solution map
+    $S(x)\mathrel{\mathop{\ordinarycolon}}$= $\operatorname*{argmin}_{z\in F_{P}(x)}~q^{\top}z,$ (25)
     is on the boundary of $F_{P}(x)$, i.e., $S(x)\subset\partial F_{P}(x)$.
   • iii)
-    For every $x$, the MFCQ holds at any $z\in F_{P}(x)$.
+    For all $x\in\mathcal{A}$, the MFCQ holds at any $z\in F_{P}(x)$.
 Proof.
-To avoid breaking the flow of the exposition, please see the note in Appendix -B. ∎
-Based on Proposition 1, one can prove that $h_{r}(x)$ can be chosen as a NBF candidate in the sense of (2).
+i) First, it is shown that $F_{P}(x)\neq\emptyset,\forall x\in\mathcal{A}$. Recalling
+$\operatorname{Int}(\Phi(x))=\left\{c\mid a_{j}^{\top}(x)c<b_{j}(x),~\forall j\in\mathcal{J}\right\},$
+for $x\in\mathcal{A}$, since $\operatorname{Int}(\Phi(x))$ is nonempty and uniformly bounded (according to Assumption 1), one can find a sufficiently small $\varepsilon_{0}(x)>0$ and a point $c_{0}(x)\in\mathcal{D}_{y}$, such that
+$a_{j}^{\top}(x)c_{0}(x)+\varepsilon_{0}(x)<b_{j}(x),~\forall j\in\mathcal{J}.$
+For every $j$, since $\|a_{j}(x)\|\neq 0$ as in Assumption 1, there exists $c_{2}(x)\neq 0$, s.t.
+$\varepsilon_{0}=a_{j}^{\top}c_{2}=\|a_{j}\|\underbrace{\|c_{2}\|\cos\langle a_{j},c_{2}\rangle}_{=\mathrel{\mathop{\ordinarycolon}}r_{j}(x)}>0.$
+Let $r_{0}\mathrel{\mathop{\ordinarycolon}}=\min_{j\in\mathcal{J}}~r_{j}$, one has $r_{0}>0$, $\|a_{j}\|r_{0}\leq\varepsilon_{0}$, and
+$a_{j}^{\top}(x)c_{0}(x)+\|a_{j}(x)\|r_{0}(x)<b_{j}(x),~\forall j\in\mathcal{J},$
+which means that $\operatorname{Int}(F_{P}(x))\neq\emptyset$.
+ii) The following shows the uniform boundedness of $F_{P}(x)$. Recalling the expression of $F_{P}(x)$ in (13):
+$F_{P}(x)\mathrel{\mathop{\ordinarycolon}}=\left\{z\mid r\geq 0,{a}_{j}^{\top}c+\|{a}_{j}\|r\leq{b}_{j},\forall j\in\mathcal{J}\right\},$
+from which one knows that
+$0\leq r\leq$ $\max_{c\in\Phi(x)}\min_{j\in\mathcal{J}}\frac{{b}_{j}(x)-{a}_{j}^{\top}(x)c}{\|{a}_{j}(x)\|}$
+$=$ $\max_{c\in\Phi(x)}d(c,\partial\Phi(x))\leq\max_{c\in\mathcal{D}_{y}}d(c,\partial\mathcal{D}_{y})=\mathrel{\mathop{\ordinarycolon}}r_{d},$
+where $d(c,\mathcal{C})\mathrel{\mathop{\ordinarycolon}}=\min_{c^{\prime}\in\mathcal{C}}\|c-c^{\prime}\|$ denotes the distance function from a point $x$ and a set $\mathcal{C}$, the compact set $\mathcal{D}_{y}$ is derived from Assumption 1 and $r_{d}$ is a positive constant.
+Furthermore, one has
+$F_{P}(x)\subset\Phi(x)\times[0,r_{d}]\subset\mathcal{D}_{y}\times[0,r_{d}],$
+which means that the set $F_{P}(x)$ is uniformly bounded, and functions $r^{*}(x)$ and $h_{r}(x)$ are also uniformly bounded.
+Due to the fact that (12) is a LP problem, the optimization value is obtained at a point $z$ only if $z$ is on the boundary $\partial F_{P}(x)$.
+iii) The MFCQ condition holds for the optimization problem in (12) since it is a LP piecewise at $x$.
+Recalling the fact in [23] that for $\mathcal{C}^{1}$ convex programs with affine inequalities, the MFCQ condition is equivalent to the Slater condition.
+Since we have proved in Theorem 1 i) that $\operatorname{Int}(F_{P}(x))\neq\emptyset,\forall x\in\mathcal{A}$, there exists a point $z\in\operatorname{Int}(F_{P}(x))$ such that the inequalities in (13) holds strictly, hence the Slater condition (as well as the MFCQ condition) naturally holds at any $z\in F_{P}(x)$.
+∎
+Based on Proposition 1, one can prove that $h_{r}(x)$ is a NBF candidate by verifying the conditions in Definition 1.
 Theorem 1.
-The Parametric LP defined implicit function $h_{r}(x)$ in (14) is locally Lipschitz continuous, and it is a NBF candidate on closed set $\mathcal{S}:=\{x:h_{r}(x)\geq 0\}$.
+The Parametric LP defined implicit function $h_{r}(x)$ in (12) is locally Lipschitz continuous on $\mathcal{A}$, and it is a NBF candidate on the closed set $\mathcal{S}\mathrel{\mathop{\ordinarycolon}}=\{x\mid h_{r}(x)\geq 0\}\subset\mathcal{A}$.
 Proof.
-Recalling Proposition 1 i) and iii) that ${F}_{P}(x)$ is uniformly bounded for all $x$ and the MFCQ condition holds, according to Lemma 2, $h_{r}(x)$ is locally Lipschitz continuous at every $x$.
-From (14), the conditions in (2) hold for $h_{r}(x)$, hence $h_{r}(x)$ is a NBF candidate on $\mathcal{S}$.
+Recalling Proposition 1 ii) and iii) that ${F}_{P}(x)$ is uniformly bounded for $x\in\mathcal{A}$ and the MFCQ condition holds, the compact set $C_{0}$ required in Lemma A.1 exists, hence the conditions in Lemma A.1 are satisfied, thus $h_{r}(x)$ is locally Lipschitz continuous.
+Since $\mathcal{S}$ is defined by $h_{r}(x)$, $h_{r}(x)$ is a NBF candidate on the set $\mathcal{S}$ according to Definition 1.
 ∎
-Before going to develop the theoretical results, the geometric structure of the solution map $S(x)$ is discussed, which is helpful to understand the expression of $h_{r}^{\prime}(x;\dot{x})$ in (28).
-First, from the definition (25), $S(x)$ can be rewritten as
-$S(x)$ $=T(x)\times\left\{r\mid r=r^{*}(x)\right\},$
-where
-$T(x):=\bigcap_{j\in\mathcal{J}}\left\{c\mid a_{j}^{\top}(x)c\leq b_{j}(x)-\|a_{j}(x)\|r^{*}(x)\right\}.$
-Since $r^{*}(x)\geq 0$, one has
-$T(x)\subseteq\left\{c\mid a_{j}^{\top}(x)c\leq b_{j}(x),\forall j\in\mathcal{J}\right\}=\Phi(x)\subset\mathcal{D}_{y}.$
-Because $T(x)$ is expressed as the intersection of halfspcaes, it is still a (convex and compact) polytope in $\mathcal{D}_{y}\subset\mathbb{R}^{l}$, whose dimension, lower than $\Phi(x)$, may be $10,1,2,...,l-$ 1.
-In the above, dimension 00 means that $T(x)$ reduces to a singleton.
 IV-B Building the NCBF Condition
-This section aims to establish the NCBF condition based on existing results about the directional differentiability of parametric convex programs in Lemma 3.
-In detail, recalling the parameteric linear program in (14):
-$v_{r}(x):=\min_{z\in F_{P}(x)}~{}q^{\top}z=-h_{r}(x)-\epsilon_{0},$
-then the directional derivatives of $v_{r}(x)$ and $h_{r}(x)$ are studied.
-First, according to the feasible set $F_{P}(\cdot)$ in (15), the active index set [^1] Inspired by [19], $\mathcal{J}_{0}({z},x)$ can be replaced by the socalled almost-active index set $\mathcal{J}_{\varepsilon}({z},x):=\{j\in\mathcal{J}\mid\|\tilde{a}_{j}(x)^{\top}{z}-b_{j}(x)\|\leq\varepsilon\}$ for real implementation, where $\varepsilon$ is a sufficiently small scalar, such as $\varepsilon=10^{-4}$.
-is defined for every ${x}$ and feasible ${z}\in F_{P}({x})$ as:
-$\mathcal{J}_{0}({z},x):=$$\left\{j\in\mathcal{J}\mid\tilde{a}_{j}(x)^{\top}{z}=b_{j}(x)\right\}$ (26)
-$=$ $\left\{j\in\mathcal{J}\mid a_{j}^{\top}(x)c+\|a_{j}(x)\|r=b_{j}(x)\right\},$
-where $\tilde{a}_{j}(x):=\operatorname{col}(a_{j}(x),\|a_{j}(x)\|)$.
-For every $\bar{z},\bar{x}$ satisfying $\bar{z}\in S(\bar{x})$, recalling $S(\bar{x})\subset\partial F_{P}(\bar{x})$, it holds that
+This section aims to establish the NCBF condition based on existing results about the directional differentiability of parametric convex programs in Lemma A.2.
+First, according to the feasible set $F_{P}(\cdot)$ in (13), the active index set [^1] Inspired by [19], $\mathcal{J}_{0}({z},x)$ can be replaced by the socalled almost-active index set $\mathcal{J}_{\varepsilon}({z},x)\mathrel{\mathop{\ordinarycolon}}=\{j\in\mathcal{J}\mid\|\tilde{a}_{j}^{\top}(x){z}-\tilde{b}_{j}(x)\|\leq\varepsilon\}$ for real implementation, where $\varepsilon$ is a sufficiently small scalar, such as $\varepsilon=10^{-4}$.
+is defined for every x{x} and feasible ${z}\in F_{P}({x})$ as:
+$\mathcal{J}_{0}({z},x)\mathrel{\mathop{\ordinarycolon}}$= $\left\{j\in[N+1]\mid\tilde{a}_{j}^{\top}(x){z}=\tilde{b}_{j}(x)\right\}.$ (26)
+For every $\bar{z},\bar{x}$ satisfying $\bar{z}\in S(\bar{x})$, recalling $S(\bar{x})\subset\partial F_{P}(\bar{x})$ in Proposition 1 iii), it holds that
 $\mathcal{J}_{0}(\bar{z},\bar{x})\neq\emptyset,\text{ i.e., }|\mathcal{J}_{0}(\bar{z},\bar{x})|\geq 1.$
-Based on (24), one can further define the active matrix of $\tilde{\mathbf{A}}(x)$ and the active vector of $\tilde{\mathbf{b}}(x)$ for every $x$ as:
-$\tilde{\mathbf{A}}_{\mathcal{J}_{0}(\bar{z},\bar{x})}(x):=\left[\begin{array}[]{c}\tilde{a}_{j_{1}}(x)^{\top}\\ \tilde{a}_{j_{2}}(x)^{\top}\\ ......\\ \tilde{a}_{j_{m}}(x)^{\top}\end{array}\right],\tilde{\mathbf{b}}_{\mathcal{J}_{0}(\bar{z},\bar{x})}(x):=\left[\begin{array}[]{c}b_{j_{1}}(x)\\ b_{j_{2}}(x)\\ ......\\ b_{j_{m}}(x)\end{array}\right],$ (27)
+Based on (24) and (26), one can further define the active matrix of $\tilde{\mathbf{A}}(x)$ and the active vector of $\tilde{\mathbf{b}}(x)$ for every $x$:
+$\tilde{\mathbf{A}}_{\mathcal{J}_{0}(\bar{z},\bar{x})}(x)\mathrel{\mathop{\ordinarycolon}}=\!\!\left[\!\!\!\begin{array}[]{c}\tilde{a}_{j_{1}}^{\top}(x)\\ \tilde{a}_{j_{2}}^{\top}(x)\\ ......\\ \tilde{a}_{j_{m}}^{\top}(x)\end{array}\!\!\!\right],\tilde{\mathbf{b}}_{\mathcal{J}_{0}(\bar{z},\bar{x})}(x)\mathrel{\mathop{\ordinarycolon}}=\!\!\left[\!\!\!\begin{array}[]{c}\tilde{b}_{j_{1}}(x)\\ \tilde{b}_{j_{2}}(x)\\ ......\\ \tilde{b}_{j_{m}}(x)\end{array}\!\!\!\right],$ (27)
 where $j_{1},j_{2},...,j_{m}$ represent all indices in $\mathcal{J}_{0}(\bar{z},\bar{x})$.
-In the above, variables “$\bar{z},\bar{x}$”, appearing in the active set $\mathcal{J}_{0}(\cdot,\cdot)$, are introduced to make a difference from the variable “$x$” in $\tilde{a}_{j}(x)$ and $b_{j}(x)$.
-Then the directional derivative of $h_{r}(x)$ along the system can be obtained through Lemma 3.
+In the above, variables $\bar{z},\bar{x}$, appearing in the active set $\mathcal{J}_{0}(\cdot,\cdot)$, are introduced to make a difference from the variable $x$ in $\tilde{a}_{j}(x)$ and $\tilde{b}_{j}(x)$.
+Then the directional derivative of $h_{r}(x)$ along the system can be obtained through Lemma A.2.
 Proposition 2.
-The directional derivative of $h_{r}(x)$ in (14) at $x$ along $\dot{x}$ can be expressed as
+The directional derivative of $h_{r}(x)$ in (12) at $x$ along $\dot{x}$ is
 $h^{\prime}_{r}(x;\dot{x})=\!\!\max_{z\in S(x)}\min_{\mu\in M(z,x)}\!\!-\partial_{x}L(\mu,z,x)\dot{x},$ (28)
 where the solution map $S(x)$ is defined as in (25),
-$\partial_{x}L(\mu,z,x)$ $=\sum_{j\in\mathcal{J}_{0}({z},{x})}\mu_{j}\left(z^{\top}\frac{\partial\tilde{{a}}_{j}}{\partial x}(x)-\frac{\partial{{b}}_{j}}{\partial x}(x)\right),$ (29)
-and the dual feasible set
+$\partial_{x}L(\mu,z,x)$ $=\!\sum_{j\in\mathcal{J}_{0}({z},{x})}\mu_{j}\left(z^{\top}\frac{\partial\tilde{{a}}_{j}}{\partial x}(x)-\frac{\partial\tilde{b}_{j}}{\partial x}(x)\right),$ (29)
+and the set of multipliers
 $M(z,x)=$ $\left\{\mu\mid\tilde{\mathbf{A}}^{\top}_{\mathcal{J}_{0}({z},{x})}(x)\mu=-q,~\mu\geq 0\right\},$ (30)
@@
 Proof.
-To avoid breaking the flow of the exposition, please see the note in Appendix -B. ∎
+This proposition is the corollary of Lemma A.2 with the boundness of set $M(z,x)$ to be proven. To prove this, first define
+$v_{r}(x)\mathrel{\mathop{\ordinarycolon}}=\min_{z\in F_{P}(x)}~q^{\top}z,$
+then $h_{r}(x)$ can be expressed as $h_{r}(x)=-v_{r}(x)-\epsilon_{0}$.
+According to (A.3), the local Lagrangian function near $\bar{x},\bar{z}$ is defined as:
+$L(\mu,z,x)=q^{\top}z+\mu^{\top}(\tilde{\mathbf{A}}_{\mathcal{J}_{0}(\bar{z},\bar{x})}(x)z-\tilde{\mathbf{b}}_{\mathcal{J}_{0}(\bar{z},\bar{x})}(x)),$
+where $\mu\in\mathbb{R}^{|\mathcal{J}_{0}({z},x)|}$, and $\tilde{\mathbf{A}}_{\mathcal{J}_{0}},\tilde{\mathbf{b}}_{\mathcal{J}_{0}}$ are defined in (27).
+Then, the partial derivative of $L(\mu,z,x)$ w.r.t.
+vectors $z$ and $x$ are respectively given by:
+$\partial_{z}L(\mu,z,x)$ $=q^{\top}+\mu^{\top}\tilde{\mathbf{A}}_{\mathcal{J}_{0}(\bar{z},\bar{x})}(x),$
+$\partial_{x}L(\mu,z,x)$ $=\mu^{\top}\left(\frac{\partial}{\partial x}\left(\tilde{\mathbf{A}}_{\mathcal{J}_{0}(\bar{z},\bar{x})}(x)z\right)-\frac{\partial}{\partial x}\tilde{\mathbf{b}}_{\mathcal{J}_{0}(\bar{z},\bar{x})}(x)\right)$
+$=\sum_{j\in\mathcal{J}_{0}(\bar{z},\bar{x})}\mu_{j}\left(z^{\top}\frac{\partial\tilde{{a}}_{j}}{\partial x}(x)-\frac{\partial\tilde{{b}}_{j}}{\partial x}(x)\right),$
+and $M(z,x)\mathrel{\mathop{\ordinarycolon}}=\left\{\mu\mid\partial_{z}L(\mu,z,x)=0^{\top},~\mu\geq 0\right\}$.
+In the above, the partial derivative of $L(\mu,z,x)$ w.r.t.
+variable $z,x$ at the point $\bar{z},\bar{x}$ are calculated.
+By further choosing $\bar{z}=z$ and $\bar{x}=x$, the expressions in (29) and (30) can be obtained.
+Since MFCQ condition holds and $F_{P}(x)$ is compact as proved in Proposition 1, then the conditions of Lemma A.2 hold, from which one has
+$h^{\prime}_{r}(x;\dot{x})$ $=-v^{\prime}_{r}(x;\dot{x})=-\min_{z\in S(x)}\max_{\mu\in M(z,x)}\partial_{x}L(\mu,z,x)\dot{x}$
+$=\max_{z\in S(x)}\min_{\mu\in M(z,x)}-\partial_{x}L(\mu,z,x)\dot{x}.$
+Hence, the expression in (28) is obtained.
+The set $M(z,x)$ in (30) can be further rewritten as:
+$M(z,x)$ $=\left\{\mu\mid\sum_{j\in\mathcal{J}_{0}({z},{x})}\tilde{a}_{j}(x)\mu_{j}=-q,~\mu\geq 0\right\}$
+$=M_{1}(z,x)\cap M_{2}(z,x),$
+where
+$M_{1}(z,x)$ $\mathrel{\mathop{\ordinarycolon}}=\left\{\mu\mid\sum_{j\in\mathcal{J}_{0}({z},{x})}a_{j}(x)\mu_{j}=0\right\},$
+$M_{2}(z,x)$ $\mathrel{\mathop{\ordinarycolon}}=\left\{\mu\mid\sum_{j\in\mathcal{J}_{0}({z},{x})}\|a_{j}(x)\|\mu_{j}=1,~\mu\geq 0\right\}.$
+From the structure of $M_{2}(z,x)$, since $\|a_{j}(x)\|>0$ from Assumption 1, $M_{2}(z,x)$ is a bounded cone [21], hence its subset $M(z,x)$ is bounded naturally. ∎
 Based on Proposition 2 and Definition 2, the NCBF condition to guaranteeing $h_{r}(x(t))\geq 0,~\forall t$ along the system (6a) can be briefly expressed as follows:
-$\max_{\begin{subarray}{c}z\in S(x)\\ u\in\Psi(x)\end{subarray}}\min_{\mu\in M(z,x)}-\partial_{x}L(\mu,z,x)\left(f+gu\right)\geq-\alpha(h_{r}(x)).$ (31)
+$\!\max_{\begin{subarray}{c}z\in S(x)\\ u\in\Psi(x)\end{subarray}}\min_{\mu\in M(z,x)}\!\!-\partial_{x}L(\mu,z,x)\left(f+gu\right)\geq-\alpha(h_{r}).$ (31)
 It is worth mentioning that condition (31) is more complex than the regular NCBF condition (5).
@@
 Theorem 2.
-The BF candidate $h_{r}(\cdot)$ is further a NCBF on $\mathcal{S}$, i.e., the NCBF condition in (4) holds for $h_{r}(\cdot)$, if there exist $\alpha,\epsilon>0$, for all $x\in\mathcal{D}_{\epsilon}$, $\exists u\in\Psi(x)$, such that
+The BF candidate $h_{r}(\cdot)$ is further a NCBF on $\mathcal{S}$, i.e., the NCBF condition in (3) holds for $h_{r}(\cdot)$, if there exist $\alpha,\bar{\epsilon}>0$, for all $x\in\mathcal{D}_{\bar{\epsilon}}\mathrel{\mathop{\ordinarycolon}}=\{x\mid h_{r}(x)\geq-\bar{\epsilon}\}$, $\exists u\in\Psi(x)$, such that
 $\min_{\mu\in M(z^{*}(x),x)}-\partial_{x}L(\mu,z^{*}(x),x)\left(f+gu\right)\geq-\alpha h_{r},$ (32)
-where $z^{*}(x)=\operatorname{col}{(c^{*}(x),r^{*}(x))}$ is chosen from $S(x)$ and the expression of $\partial_{x}L(\cdot)$ is given in (29).
+where $z^{*}(x)=\operatorname{col}{(c^{*}(x),r^{*}(x))}$ is any point chosen from the set $S(x)$ and the expression of $\partial_{x}L(\cdot)$ is given by (29).
 Proof.
@@
 $=\max_{z\in S(x)}\min_{\mu\in M(z,x)}-\partial_{x}L(\mu,z,x)\left(f+gu\right)$
-$\geq~{}\min_{\mu\in M(z^{*}(x),x)}-\partial_{x}L(\mu,z^{*}(x),x)\left(f+gu\right)$
-$\geq-\alpha h_{r},$
+$\geq~\min_{\mu\in M(z^{*}(x),x)}-\partial_{x}L(\mu,z^{*}(x),x)\left(f+gu\right)\geq-\alpha h_{r},$
 which indicates the NCBF condition (31) holds. ∎
@@
 of $M(z^{*}(x),x)$ in (30) automatically, such as the open-source toolbox $\mathrm{pypoman}$ [25]., such that
-$M(z^{*}(x),x)=\operatorname{co}\left(\{{\mu}^{k}(x)\mid k\in\mathcal{I}_{V_{M}}(x)\}\right),$ (33)
+$M(z^{*}(x),x)=\operatorname{co}\left(V_{M}(x)\right),$ (33)
 where $\operatorname{co}(\cdot)$ denotes the convex hull operator of a set, and $\mathcal{I}_{V_{M}}(x)$ is the index set of ${V_{M}}(x)$.
@@
 Corollary 1.
-The BF candidate $h_{r}(\cdot)$ is further a NCBF on $\mathcal{S}$, i.e., the NCBF condition in (4) holds for $h_{r}(\cdot)$, if there exist $\alpha,\epsilon>0$, for all $x\in\mathcal{D}_{\epsilon}$, $\exists u\in\Psi(x)$, such that
+The BF candidate $h_{r}(\cdot)$ is further a NCBF on $\mathcal{S}$, if there exist $\alpha,\bar{\epsilon}>0$, for all $x\in\mathcal{D}_{\bar{\epsilon}}\mathrel{\mathop{\ordinarycolon}}=\{x\mid h_{r}(x)\geq-\bar{\epsilon}\}$, $\exists u\in\Psi(x)$, such that
 $\Gamma(\mu^{k}(x),x,u)\geq-\alpha h_{r}(x),~\forall k\in\mathcal{I}_{V_{M}}(x),$ (34)
-where $\Gamma(\mu,x,u):=-\partial_{x}L(\mu,z^{*}(x),x)\left(f+gu\right)$ for any $\mu$.
+where $\Gamma(\mu,x,u)\mathrel{\mathop{\ordinarycolon}}=-\partial_{x}L(\mu,z^{*}(x),x)\left(f+gu\right)$ for any $\mu$.
 Proof.
 From the expression of $\partial_{x}L(\cdot)$ in (29), one can know that the scalar function $\partial_{x}L(\mu,z^{*},x)\left(f+gu\right)$ is a linear function of $\mu,z^{*},u$ respectively, hence the minimization optimization about $\mu$ in (32) is a LP at every $x$.
-Recalling the fact in [25, Sec 2.5, Theorem and Corollary 2] that the minimum value of a LP is obtained at the vertexes of the polytope, then
-$\min_{\mu\in M(z^{*}(x),x)}\Gamma(\mu,x,u)$ $=\min_{\mu\in V_{M}(x)}\Gamma(\mu,x,u)$
-$=\min_{{k}\in\mathcal{I}_{V_{M}}(x)}\Gamma(\mu^{k},x,u),$
+Recalling the fact in [26, Sec 2.5, Theorem and Corollary 2] that the minimum value of a LP is obtained at the vertexes of the polytope, then
+$\min_{\mu\in M(z^{*}(x),x)}\Gamma(\mu,x,u)$ $=\min_{{k}\in\mathcal{I}_{V_{M}}(x)}\Gamma(\mu^{k},x,u),$
 which can be further combined with (32) to verify (34). ∎
@@
 $\bar{u}^{*}(x)=\operatorname*{argmin}_{u,\delta}~(u-u_{0})^{\top}Q(u-u_{0})+\gamma\delta^{2},$ (35a)
-[…]
+$\text{s.t. }\Gamma(\mu^{k}(x),x,u)\geq-\alpha h_{r}(x)-\delta,~\forall k\in\mathcal{I}_{V_{M}}(x),$ (35b)
 $\quad\quad u\in\Psi(x),~\delta\geq 0,$ (35c)
-where $\bar{u}^{*}(x):=\operatorname{col}(u^{*}(x),\delta^{*}(x))$, $u_{0}$ is any given continuous nominal control law, and the slack variable $\delta$ is introduced to guarantee the feasibility of such QP.
-$Q\in\mathbb{R}^{m\times m},\gamma>0$ are the (positive-definite) weighting matrix and scalar.
+where $\bar{u}^{*}(x)\mathrel{\mathop{\ordinarycolon}}=\operatorname{col}(u^{*}(x),\delta^{*}(x))$, $u_{0}$ is any given continuous nominal control law, and the slack variable $\delta$ is introduced to guarantee the feasibility of such QP.
+$Q\in\mathbb{R}^{m\times m},\gamma>0$ are the (positive-definite) weighting matrix and scalar respectively, where the penalty scalar $\gamma$ is much larger than the maximal eigenvalue of $Q$.
 The structure of the proposed safety filter (35) is summarized in Fig.
-2.
-Figure 2: A summary of the structure of the proposed safety filter (35).
+1 (b).
 Till now, the final theorem is shown to conclude this paper.
 Theorem 3.
-Consider the designed safety filter in the form of (35). For any system trajectory $x(t)$ starting from $x_{0}\in\mathcal{S}=\{x:h_{r}(x)\geq 0\}$, if
-$u^{*}(x(t))\text{ is continuous, and }\delta^{*}(x(t))\equiv 0,$ (36)
-then $x(t)\in\mathcal{S},\forall t\geq 0$, which implies that the PVM problem is solved. As a consequence,
-$\Phi(x(t))\neq\emptyset,~{}u^{*}(x(t))\in\Psi(x(t)),\quad\forall t\geq 0,$ (37)
-i.e., the output set-valued mapping never be empty and meanwhile the input constraint is never violated.
+If for the designed control law $\bar{u}^{*}(x)=\operatorname{col}(u^{*}(x),\delta^{*}(x))$ in (35), it holds at every $x\in\mathcal{S}=\{x\mid h_{r}(x)\geq 0\}$ that ${u}^{*}(x)$ is locally bounded and measurable, and the slack variable $\delta^{*}(x)$ is bounded by
+$\delta^{*}(x)\leq\alpha\epsilon_{0}.$ (36)
+Then the set $\mathcal{S}$ is forward invariant. As a consequence, for any trajectory $x(t)$ starting from $x_{0}\in\mathcal{S}$, it holds that
+$x(t)\in\mathcal{A},~u^{*}(x(t))\in\Psi(x(t)),\quad\forall t\geq 0,$
+i.e., the output set is always nonempty and input constraint is never violated.
 Proof.
 Since Eq.
-(35b) denotes the soft constraints, the feasibility of the QP-based safety filter is guaranteed.
-The solution of closed-loop system is well-defined by the assumption that $u^{*}(x(t))$ is continuous.
+(35b) denotes the soft constraints, the feasibility of the QP-based safety filter is guaranteed whenever $\Psi(x)\neq\emptyset$.
+The solution of closed-loop system is well-defined by the assumption that ${u}^{*}(x(t))$ is be measurable and locally bounded [19].
 Meanwhile, the input constraint is never violated due to the hard constraints in (35c).
 Under the designed safety filter (35), one has
-$h^{\prime}_{r}(x(t);\dot{x}(t))\geq-\alpha h_{r}(x(t))-\delta^{*}(x(t)),$ (38)
-where $\dot{x}(t)=f_{0}(x(t),u^{*}(x(t)))$.
-When $\delta^{*}(x(t))\equiv 0$, the analysis is obvious.
-In fact, from (38), $h^{\prime}_{r}(x(t);\dot{x}(t))\geq-\alpha h_{r}(x(t))$, by referring to Lemma 1, $h_{r}(x(t))\geq 0$, i.e., $x(t)\in\mathcal{S},\forall t\geq 0$.
-Recalling (12), the volume of $\Phi(x(t))$ is always greater than a threshold $\epsilon$.
+$(r^{*})^{\prime}(x(t);\dot{x}(t))$ $=h^{\prime}_{r}(x(t);\dot{x}(t))$ (37)
+$\geq-\alpha h_{r}(x(t))-\delta^{*}(x(t))$
+$=-\alpha r^{*}(x(t))+\alpha\epsilon_{0}-\delta^{*}(x(t)).$
+When the inequality (36) holds, (37) will be reduced to $(r^{*})^{\prime}(x(t);\dot{x}(t))\geq-\alpha r^{*}(x(t))$.
+By referring to Lemma 1, the set $\mathcal{S}$ is forward invariant.
+Recalling the relationship in (11), one has $x(t)\in\mathcal{A}=\{x\mid\operatorname{Int}(\Phi(x))\neq\emptyset\},~\forall t\geq 0$.
 ∎
 Remark 4.
-In Theorem 3, two conditions in (36) are attached for the sake of completeness. They could be further relaxed. i) Known from (14), the inequality in (38) can be equivalently rewritten as:
-${r^{*}}^{\prime}(x(t);\dot{x}(t))\geq-\alpha{r^{*}}(x(t))+{\alpha\epsilon_{0}-\delta^{*}(x(t))}.$ (39)
-This reveals that if
-$\|\delta^{*}(x(t))\|_{\mathcal{L}_{\infty}}:=\sup_{t}\|\delta^{*}(x(t)\|<\alpha\epsilon_{0},$
-then ${r^{*}}^{\prime}(x(t);\dot{x}(t))>-\alpha{r^{*}}(x(t))$, which means that
-${r^{*}}(x(t))>0,\forall t\geq 0.$
 […]
@@
 $\dot{p}_{x}=v\cos\theta,~\dot{p}_{y}=v\sin\theta,~\dot{v}=a,~\dot{\theta}=\omega.$
-Variables $x=[{p}_{x},{p}_{y},v,\theta]^{\top}$ are states, and $u=[a,\omega]^{\top}$ are control inputs, restricted in $\mathcal{U}=[0,2]\times[0,2]$.
+Variables $x=[{p}_{x},{p}_{y},v,\theta]^{\top}$ are states, and $u=[a,\omega]^{\top}$ are control inputs, restricted in $\mathcal{U}=[-2,2]\times[-2,2]$.
 The reference controller in [9], with parameters $k_{p},k_{\theta},k_{v}$, is applied as the nominal control $u_{0}$ to drive the vehicle from the initial point to the goal.
@@
 $h_{i}(x)=\|[{p}_{x},{p}_{y}]^{\top}-[{o}_{xi},{o}_{yi}]^{\top}\|^{2}-R^{2},$
-where $[{o}_{xi},{o}_{yi}]^{\top}$ is the center of the $i$-th obstacle and $R$ is the pre-defined safety distance.
+where ${o}_{xi},{o}_{yi}]^{\top}$ is the center of the $i$-th obstacle and $R$ is the pre-defined safety distance.
 The positions of obstacles, initial and goal points can be seen from Fig.
-3.
+2.
 The HOCBF [28] (with order $2$ and parameters $\alpha_{1}=10.0,\alpha_{2}=6.0$) is utilized to transform the position constraints $h_{i}(x)\geq 0$ to the affine constraints on input $u$, denoted as $u\in\mathcal{U}_{c}(x)$.
@@
 (b) The time response of $r^{*}(x(t))$.
-(c) The time responses of $\delta^{*}(x(t))$ and $u^{*}(x(t))$.
+(c) Time responses of $\delta^{*}(x(t))$ and $u^{*}(x(t))$.
 Figure 2: Simulation results in different scenarios: Case 1 ($k_{v}=0.5$), Case 2 ($k_{v}=1.0$), and Case 3 ($k_{v}=2.0$).
@@
 To evaluate the performance sufficiently, the nominal control law $u_{0}$ with three different values of the gain $k_{v}$ are introduced in the simulation so as to produce different moving behavior, named as Case 1 ($k_{v}=0.5$), Case 2 ($k_{v}=1.0$), and Case 3 ($k_{v}=2.0$), as seen in Fig.
-3.
+2.
 Fig.
-3(a) shows the trajectories of the vehicle in the $XY$ plane.
+2(a) shows the trajectories of the vehicle in the $XY$ plane.
 It can be observed that under the proposed safety filter (35), the vehicle successfully avoids the obstacles and reaches the goal in all cases.
-Under the method being compared (40), in Case 2, the vehicle can successfully avoid the obstacles and reaches the goal, while in Case 1 and Case 3, the vehicle would get stuck in the obstacles.
+Under the compared method (38), in Case 2, the vehicle can successfully avoid the obstacles and reaches the goal, while in Case 1 and Case 3, the vehicle would get stuck in the obstacles.
 Fig.
-3(b) illustrates the time response of $r^{*}(x(t))$, which demonstrates the range of the feasible space of these two different safety filters.
+2(b) illustrates the time response of $r^{*}(x(t))$, which demonstrates the range of the feasible space of these two different safety filters.
 From it, one can observe that $r^{*}(x(t))$ is always greater than zero and greater than $\epsilon_{0}=0.6$ in most cases, which shows that the proposed method can maintain feasibility.
@@
 In Fig.
-3(c), the time responses of slack variable $\delta^{*}(x(t))$ and inputs $u^{*}(x(t))=[a(x(t)),\omega(x(t))]^{\top}$ under the proposed safety filter (35) in the above three cases are shown.
-One can observe that in all cases, $\delta^{*}(x(t))$ is equal to 00 approximately and much lower than the threshold $\alpha\epsilon_{0}=9$.
+2(c), the time responses of slack variable $\delta^{*}(x(t))$ and inputs $u^{*}(x(t))=[a(x(t)),\omega(x(t))]^{\top}$ under the proposed safety filter (35) in the above three cases are shown.
+One can observe that in all cases, $\delta^{*}(x(t))$ is equal to 0 approximately and much lower than the threshold $\alpha\epsilon_{0}=9$.
 Besides, the inputs $u^{*}(x(t))$ are continuous in all cases.
@@
 The biggest limitation of this work is that the continuity of the proposed safety filter is not discussed formally.
-Future work involves designing continuous safety filters to tackle PVM problems, extending the proposed analysis and synthesis framework to other relevant problems and so on.
--A Nonsmooth Parametric Optimization
-Let $\mathbb{T}\subset\mathbb{R}^{n}$ denote some open parameter set. For all $\bar{x}\in\mathbb{T}$, consider the convex parametric program with only inequalities:
-$v(\bar{x}):=\min_{\bar{z}}$ $~{}\tilde{f}(\bar{z},\bar{x}),$ (A.1)
-s.t. $\bar{z}\in{F}(\bar{x}):=\left\{\bar{z}\in\mathbb{R}^{n_{z}}\mid\tilde{g}_{j}(\bar{z},\bar{x})\leq 0,j\in\mathcal{J}\right\},$
-with $\mathcal{J}:=\{1,2,\ldots,N\}$. For convenience, let the set of (global) minimizers according to (A.1) be denoted as
-$S(\bar{x}):=\left\{\bar{z}\in F(\bar{x})\mid\tilde{f}(\bar{z},\bar{x})=v(\bar{x})\right\}.$ (A.2)
-For $\bar{x}\in\mathbb{T}$ and feasible $\bar{z}\in F(\bar{x})$, the active index set is defined as $\mathcal{J}_{0}(\bar{z},\bar{x}):=\left\{j\in\mathcal{J}\mid\tilde{g}_{j}(\bar{z},\bar{x})=0\right\}$.
+Future works involve designing continuous and smooth safety filters to tackle PVM problems, extending the proposed analysis and synthesis framework to other relevant problems and so on.
+Appendix A Appendix
+A-A Nonsmooth Parametric Optimization
+Let $\mathbb{T}\subset\mathbb{R}^{n}$ denote some open parameter set. For all $\bar{x}\in\mathbb{T}$, consider the convex parametric program:
+$v(\bar{x})\mathrel{\mathop{\ordinarycolon}}=\min_{\bar{z}}$ $~\tilde{f}(\bar{z},\bar{x}),$ (A.1)
+s.t. $\bar{z}\in{F}(\bar{x})\mathrel{\mathop{\ordinarycolon}}=\left\{\bar{z}\in\mathbb{R}^{n_{z}}\mid\tilde{g}_{j}(\bar{z},\bar{x})\leq 0,j\in\mathcal{J}\right\},$
+with $\mathcal{J}\mathrel{\mathop{\ordinarycolon}}=\{1,2,\ldots,N\}$. For convenience, let the set of (global) minimizers according to (A.1) be denoted as
+$S(\bar{x})\mathrel{\mathop{\ordinarycolon}}=\left\{\bar{z}\in F(\bar{x})\mid\tilde{f}(\bar{z},\bar{x})=v(\bar{x})\right\}.$ (A.2)
+For $\bar{x}\in\mathbb{T}$ and feasible $\bar{z}\in F(\bar{x})$, the active index set is defined as $\mathcal{J}_{0}(\bar{z},\bar{x})\mathrel{\mathop{\ordinarycolon}}=\left\{j\in\mathcal{J}\mid\tilde{g}_{j}(\bar{z},\bar{x})=0\right\}$.
 Throughout this section, $\tilde{f},\tilde{g}_{j},j\in\mathcal{J}$, are continuously differentiable on $\mathbb{R}^{n_{z}}\times\mathbb{T}$, notation $\tilde{f},\tilde{g}_{j}\in\mathcal{C}^{1}$.
-The following Constraint Qualification conditions can be found in [18, (1.3) and (1.4)] and [27, (11) and (12)].
-Definition 3.
+The following Constraint Qualification conditions can be found in [18, (1.3) and (1.4)] and [23, (11) and (12)].
+Definition A.1.
 There are two constraint qualification conditions, respectively defined as:
@@
 The continuity and directional differentiability properties of the value function of parametric programs with only inequalities have been studied in [18].
-Lemma 2.
+Lemma A.1.
 [18, Lemma 6.2] Consider the parametric convex program in (A.1) and assume there exists a neighborhood $B_{\varepsilon_{0}}(\bar{x}),\varepsilon_{0}>0$, such that $\cup_{x\in B_{\varepsilon_{0}}(\bar{x})}F(x)\subset C_{0}$, with $C_{0}$ compact.
@@
 $L(\mu,z,x)=\tilde{f}(z,x)+\sum_{j\in\mathcal{J}_{0}(\bar{z},\bar{x})}\mu_{j}\tilde{g}_{j}(z,x),$ (A.3)
-where ${\mu}:=\left({\mu}_{j},j\in\mathcal{J}_{0}(\bar{z},\bar{x})\right)\in\mathbb{R}^{\left|\mathcal{J}_{0}(\bar{z},\bar{x})\right|}$, collecting all multiplers $\mu_{j}$ corresponding to active constraints in $\mathcal{J}_{0}(\bar{z},\bar{x})$.
+where ${\mu}\mathrel{\mathop{\ordinarycolon}}=\left({\mu}_{j},j\in\mathcal{J}_{0}(\bar{z},\bar{x})\right)\in\mathbb{R}^{\left|\mathcal{J}_{0}(\bar{z},\bar{x})\right|}$, collecting all multiplers $\mu_{j}$ corresponding to active constraints in $\mathcal{J}_{0}(\bar{z},\bar{x})$.
 Then the partial derivatives of $L(\cdot)$ are respectively given by
-$\partial_{x}L(\mu,z,x):=$ $\partial_{x}\tilde{f}(z,x)+\sum_{j\in\mathcal{J}_{0}(\bar{z},\bar{x})}\mu_{j}\partial_{x}\tilde{g}_{j}(z,x),$
-$\partial_{z}L(\mu,z,x):=$ $\partial_{z}\tilde{f}(z,x)+\sum_{j\in\mathcal{J}_{0}(\bar{z},\bar{x})}\mu_{j}\partial_{z}\tilde{g}_{j}(z,x).$
-Lemma 3.
+$\partial_{x}L(\mu,z,x)\mathrel{\mathop{\ordinarycolon}}$= $\partial_{x}\tilde{f}(z,x)+\sum_{j\in\mathcal{J}_{0}(\bar{z},\bar{x})}\mu_{j}\partial_{x}\tilde{g}_{j}(z,x),$
+$\partial_{z}L(\mu,z,x)\mathrel{\mathop{\ordinarycolon}}$= $\partial_{z}\tilde{f}(z,x)+\sum_{j\in\mathcal{J}_{0}(\bar{z},\bar{x})}\mu_{j}\partial_{z}\tilde{g}_{j}(z,x).$
+Lemma A.2.
 [18, Theorem 7.3] Consider the parametric convex program in (A.1) and let $F(\bar{x})$ be compact.
-If the MFCQ condition holds at $(\bar{z},\bar{x})$ satisfying $\bar{z}\in F(\bar{x})$, then $v(\cdot)$ is directional differentiable at $\bar{x}$ in any direction $d,\|d\|=1$ with
+If MFCQ holds at $(\bar{z},\bar{x})$ satisfying $\bar{z}\in F(\bar{x})$, then $v(\cdot)$ is directional differentiable at $\bar{x}$ in any direction $d,\|d\|=1$ with
 $v^{\prime}(\bar{x};d)=\min_{\bar{z}\in S(\bar{x})}\max_{\mu\in M(\bar{z},\bar{x})}\partial_{x}L(\mu,\bar{z},\bar{x})d,$ (A.5)
-where the minimizer set $S(\cdot)$ is defined in (A.2), and
-$M(\bar{z},\bar{x})$ :=$:=\left\{\mu\mid\partial_{z}L(\mu,\bar{z},\bar{x})=0^{\top},\mu\geq 0\right\}.$ (A.6)
-If LICQ holds, then $M(\bar{z},\bar{x})$ will be a singleton, i.e., the multiplier ${\mu}$ will be uniquely determined.
-The min-max expression of directional derivative in (A.5) has also been pointed out by other literature [28, Theorem 2.4], [29, Theorem 4.24] yet under different conditions.
--B Proofs
-Proposition 1 can be proved by some standard results in LP theory. Proposition 2 is the corollary of Lemma 3 with the boundness of set $M(z,x)$ to be proven.
-Proof of Proposition 1: i) First, it is shown that $F_{P}(x)\neq\emptyset$ for all $x$. Recalling
-$\Phi(x)=\{c\mid a_{j}^{\top}(x)c\leq b_{j}(x),~{}\forall j\in\mathcal{J}\},$
-is nonempty, compact and convex, one can find a sufficiently small scalar $\varepsilon_{0}(x)>0$ and a point $c_{0}(x)$, such that
-$a_{j}^{\top}(x)c_{0}(x)+\varepsilon_{0}(x)\leq b_{j}(x),~{}\forall j=1,2,\cdots,N.$
-For every $j$, since $\|a_{j}(x)\|\neq 0$, there exists $c_{2}(x)\neq 0$, s.t.
-$\varepsilon_{0}=a_{j}^{\top}c_{2}=\|a_{j}\|\underbrace{\|c_{2}\|\cos\langle a_{j},c_{2}\rangle}_{=:r_{j}(x)}>0.$
-Defining $r_{0}:=\min_{j\in\mathcal{J}}~{}r_{j}$, one has $r_{0}>0$ and $\|a_{j}\|r_{0}\leq\varepsilon_{0}$, which leads to
-$a_{j}^{\top}(x)c_{0}(x)+\|a_{j}(x)\|r_{0}(x)\leq b_{j}(x),~{}\forall j=1,2,\cdots,N,$
-i.e., $F_{P}(x)\neq\emptyset$. Besides, $\Phi(x)$ is uniformly bounded for all $x$ as in Assumption 1, thus $F_{P}(x)$ is also uniformly bounded.
-ii) Since the feasible set $F_{P}(x)$ is uniformly bounded, $r^{*}(x)$ and $h_{r}(x)$ are also uniformly bounded.
-Due to the fact that (14) is a LP problem, the optimization value is obtained at a point $z$ only if $z$ is on the boundary $\partial F_{P}(x)$, hence $S(x)\subset\partial F_{P}(x)$ is also uniformly bounded.
-[…]
-Proof of Proposition 2: According to (A.3), the local Lagrangian function near $\bar{x},\bar{z}$ is defined as:
-$L(\mu,z,x)=q^{\top}z+\mu^{\top}(\tilde{\mathbf{A}}_{\mathcal{J}_{0}(\bar{z},\bar{x})}(x)z-\tilde{\mathbf{b}}_{\mathcal{J}_{0}(\bar{z},\bar{x})}(x)),$
-where $\mu\in\mathbb{R}^{|\mathcal{J}_{0}({z},x)|}$, and $\tilde{\mathbf{A}}_{\mathcal{J}_{0}},\tilde{\mathbf{b}}_{\mathcal{J}_{0}}$ are defined in (27).
-Then, the partial derivative of $L(\mu,z,x)$ w.r.t.
-vectors $z$ and $x$ are respectively given by:
-$\partial_{z}L(\mu,z,x)$ $=q^{\top}+\mu^{\top}\tilde{\mathbf{A}}_{\mathcal{J}_{0}(\bar{z},\bar{x})}(x),$
-$\partial_{x}L(\mu,z,x)$ $=\mu^{\top}\left(\frac{\partial}{\partial x}\left(\tilde{\mathbf{A}}_{\mathcal{J}_{0}(\bar{z},\bar{x})}(x)z\right)-\frac{\partial}{\partial x}\tilde{\mathbf{b}}_{\mathcal{J}_{0}(\bar{z},\bar{x})}(x)\right)$
-$=\sum_{j\in\mathcal{J}_{0}(\bar{z},\bar{x})}\mu_{j}\left(z^{\top}\frac{\partial\tilde{{a}}_{j}}{\partial x}(x)-\frac{\partial{{b}}_{j}}{\partial x}(x)\right),$
-and $M(z,x):=\left\{\mu\mid\partial_{z}L(\mu,z,x)=0^{\top},~{}\mu\geq 0\right\}.$ In the above, the partial derivative of $L(\mu,z,x)$ w.r.t.
-variable $z,x$ at the point $\bar{z},\bar{x}$ are calculated.
-By further choosing $\bar{z}=z,\bar{x}=x$, the expressions in (29) and (30) can be obtained.
-Since MFCQ condition holds and $F_{P}(x)$ is compact as proved in Proposition 1, then the conditions of Lemma 3 hold, from which one has
-$h^{\prime}_{r}(x;\dot{x})$ $=-v^{\prime}_{r}(x;\dot{x})$
-$=-\min_{z\in S(x)}\max_{\mu\in M(z,x)}\partial_{x}L(\mu,z,x)\dot{x}$
-$=\max_{z\in S(x)}\min_{\mu\in M(z,x)}-\partial_{x}L(\mu,z,x)\dot{x}.$
-Hence, the expression in (28) is obtained.
-The set $M(z,x)$ in (30) can be further rewritten as:
-$M(z,x)$ =$=\left\{\mu\mid\sum_{j\in\mathcal{J}_{0}({z},{x})}\tilde{a}_{j}(x)\mu_{j}=-q,~{}\mu\geq 0\right\}$
-$=M_{1}(z,x)\cap M_{2}(z,x),$
-where
-$M_{1}(z,x)$ :=$:=\left\{\mu\mid\sum_{j\in\mathcal{J}_{0}({z},{x})}a_{j}(x)\mu_{j}=0\right\},$
-$M_{2}(z,x)$ :=$:=\left\{\mu\mid\sum_{j\in\mathcal{J}_{0}({z},{x})}\|a_{j}(x)\|\mu_{j}=1,~{}\mu\geq 0\right\}.$
-[…]
+where the minimizer set $S(\cdot)$ is defined in (A.2), and the set of multipliers $M(\cdot)$ is defined as
+$M(\bar{z},\bar{x})$ $\mathrel{\mathop{\ordinarycolon}}=\left\{\mu\mid\partial_{z}L(\mu,\bar{z},\bar{x})=0^{\top},\mu\geq 0\right\}.$ (A.6)
+If LICQ holds, then $M(\bar{z},\bar{x})$ will be a singleton, i.e., the multiplier $\mu$} will be uniquely determined [18, EX.
+2.6].
+The min-max expression of directional derivative in (A.5) has also been pointed out by other literature [29, Theorem 2.4], [30, Theorem 4.24] yet under different conditions.